### v1

# The Boundaries of Tractability in Hierarchical Task Network Planning

Abstract
We study the complexity-theoretic boundaries of tractability for three classical problems in the context of Hierarchical Task Network Planning: the validation of a provided plan, whether an executable plan exists, and whether a given state can be reached by some plan. We show that all three problems can be solved in polynomial time on primitive task networks of constant partial order width (and a generalization thereof), whereas for the latter two problems this holds only under a provably necessary restriction to the state space. Next, we obtain an algorithmic meta-theorem along with corresponding lower bounds to identify tight conditions under which general polynomial-time solvability results can be lifted from primitive to general task networks. Finally, we enrich our investigation by analyzing the parameterized complexity of the three considered problems, and show that (1) fixed-parameter tractability for all three problems can be achieved by replacing the partial order width with the vertex cover number of the network as the parameter, and (2) other classical graph-theoretic parameters of the network (including treewidth, treedepth, and the aforementioned partial order width) do not yield fixed-parameter tractability for any of the three problems.
1 Introduction
Automated Planning is a core research topic whose goal is to devise computational methods capable of finding solutions to prominent planning tasks. In this work, we consider automated planning in the context of Hierarchical Task Networks (HTNs), which have received significant attention in the artificial intelligence research community. HTNs are capable of incorporating the compound, hierarchical structure of real-world tasks, as opposed to the fine-grained view of classical planning that operates directly on the state space of the system under consideration, and are one of the best-established planning formalisms in the literature Tsuneto et al. (1996); Kuter et al. (2005); Li et al. (2009); Hogg et al. (2009); Parunak et al. (2009); Sohrabi et al. (2009); Geier and Bercher (2011); Alford et al. (2015); Xiao et al. (2017); Höller et al. (2019); Behnke et al. (2019); Olz et al. (2021); Lin and Bercher (2023).
On a high level, [^1] Formal definitions are provided in the Preliminaries. HTNs can be informally described and illustrated as follows: at the top level of the task hierarchy is a sequence of tasks that is to be executed in a (partially specified) order. For instance, the task of going to work may be decomposed into three subtasks: leaving the house, going the actual distance, and getting set up at the office. Now, each of these tasks can in turn be decomposed into a set of smaller tasks, like climbing the stairs or taking a bus, which again can be decomposed, and so forth. This continues until we reach an elementary level of tasks that are no longer decomposable, which correspond to single executable actions.
The crucial feature of the model is that it separates the user-facing description of a task (what we want to do) from its internal implementation (how to do it). For instance, leaving the house always needs to be done before going to work, regardless of how it is done. However, in an HTN, we may have several options of how to decompose this task: the agent can, e.g., either climb the stairs or take the elevator. Similarly, the agent might opt for walking to work instead of going by bus.
In the literature, HTN planning is usually considered through the lens of heuristic approaches, which are evaluated against real-world problems. On the other hand, much less is known about the complexity-theoretic foundations of planning problems associated with HTNs. This lack of understanding contrasts with the major advances that have been made towards obtaining a thorough comprehension of the complexity of other fundamental problems arising in the context of artificial intelligence, such as Bayesian Network Learning Ordyniak and Szeider (2013); Grüttemeier et al. (2021a, b); Grüttemeier and Komusiewicz (2022), Clustering Ganian et al. (2020b, 2022); Kellerhals et al. (2023); Bandyapadhyay et al. (2023), Resource Allocation Bliem et al. (2016); Deligkas et al. (2022); Eiben et al. (2023), and Scheduling Ganian et al. (2020a); Bentert et al. (2023); Heeger et al. (2023). The aim of this article is to fill this gap by identifying the precise boundaries of tractability for the central computational tasks that arise in the context of HTN planning.
Considered Problems.
Perhaps the most natural algorithmic question associated with HTNs is the following: given an HTN, an initial state, and a plan, is the plan executable on the HTN? This question has been extensively investigated in the literature, where authors typically consider a plan to be formalized as either a sequence of tasks Behnke et al. (2015); Barták et al. (2021), or actions Höller et al. (2022); Lin and Bercher (2023). Here, we adopt the latter approach and note that the complexity of this problem—denoted Plan Validation—has been studied, e.g., in the context of post-optimization and repairing of solutions to HTN problems  Lin and Bercher (2023).
The second problem we consider is Plan Existence: given an HTN and some initial state, is there a way to execute all the tasks in the network? The complexity of this fundamental problem has been studied in several variants of HTN planning problems Erol et al. (1996); Geier and Bercher (2011); Alford et al. (2015); Chen and Bercher (2021).
Third, we consider the question of whether a specified target state can be reached from an initial state of the system, formalized as the State Reachability problem. This is the typical goal in classical planning Fikes and Nilsson (1971); Chapman (1987); Bäckström and Nebel (1995), so it should come as no surprise that HTN planning problems with specified target states have attracted interest as well Höller and Bercher (2021); Olz et al. (2021).
Motivated by the above considerations of states and tasks, we also define the corresponding problem of constructing a plan executing (at least) a specified set of actions. However, this Action Executability problem is in some sense a generalization of Plan Existence, and, in fact, all results that we prove for one of the problems also hold for the other. Thus, we omit Action Executability from the discussion of our results below.
Our Contribution.
The lack of results on the computational complexity of Plan Validation, Plan Existence, and State Reachability may at least partially stem from the fact that they remain computationally intractable (in particular, NP-hard) even in severely restricted settings. While the emerging intractability might seem difficult to navigate, we identify interesting and natural special cases where pockets of polynomial-time solvability show themselves. Moreover, we prove that these pockets give rise to a surprisingly rich complexity-theoretic landscape.
To be more precise, we embark by considering primitive task networks, i.e., HTNs in which there are no compound tasks. While this may seem restrictive, it is a natural place to start from a complexity-theoretic perspective: all three of the considered problems remain NP-hard on primitive task networks. In fact, as our first results, we establish the NP-hardness of Plan Validation, Plan Existence, and State Reachability even on networks which consist of collections of stars (Theorems 5 and 2). Moreover, unlike Plan Validation, we show that the latter two problems remain NP-hard even on trivial (in particular, edgeless) networks when the state space is not of fixed size (Theorem 1); this means that in order to avoid immediate intractability, we henceforth consider Plan Existence and State Reachability exclusively in the setting of constant-sized state spaces. This requirement matches analogous restrictions that were used in other planning contexts Kronegger et al. (2013); Bäckström et al. (2015).
While these initial results may seem discouraging, we prove that all three problems are polynomial-time solvable on networks which are either chains or antichains (Observations 6 and 7, Theorems 10 and 8). These tractable cases hint at cracks in the wall of computational intractability. We pry these cracks open by showing that these two restrictions actually represent special cases in a spectrum of tractable special cases: Plan Validation, Plan Existence, and State Reachability are polynomial-time solvable on networks of bounded “generalized partial order width”, that is, the partial order width of the network when ignoring isolated elements (Theorems 11, 13 and 14). This non-trivial result, established via a combination of branching techniques and deep insights into the behavior of optimal plans in the state space, unifies and generalizes the aforementioned tractability results on chains and antichains.
Building on the above results for the primitive case, in the second part of the paper, we establish an algorithmic meta-theorem that allows us to precisely describe the circumstances under which general tractability results for primitive task networks can be lifted to the compound case. Essentially, if
  1. 1.
    the total number of compound tasks in the network,
  2. 2.
    the maximum size of a network that a task can be decomposed into, and
  3. 3.
    the maximum depth of a decomposition
are all bounded by some constant, then any polynomial-time solvability result for any “reasonable” HTN planning problem on primitive task networks can be lifted to compound networks (Theorem 16). Crucially, we prove that this result is tight in the sense that tractability cannot be guaranteed if any of the three restrictions are dropped (Theorems 17, 19, 20 and 18).
In the third and final part of the paper, we push beyond classical complexity and analyze Plan Validation, Plan Existence, and State Reachability from the perspective of the more fine-grained parameterized complexity paradigm Downey and Fellows (2013); Cygan et al. (2015). In parameterized complexity, one considers the tractability of problems not only w.r.t. the input size $n$, but also w.r.t. a specified numerical parameter $k$, where the aim is to obtain algorithms which run in time at most $f(k)\cdot n^{\mathcal{O}(1)}$ (so-called fixed-parameter algorithms); problems admitting such algorithms are called fixed-parameter tractable.
The natural question from the parameterized perspective is whether the aforementioned tractability results for HTNs with constant-sized structural measures (Theorems 11, 13 and 14) can be lifted to fixed-parameter tractability when parameterized by the same measures. While analogous questions have been thoroughly explored—and often answered in the affirmative—in many other areas of artificial intelligence research Ordyniak and Szeider (2013); Eiben et al. (2021); Ganian and Korchemna (2021); Froese et al. (2022); Heeger et al. (2023), here we prove that the generalized partial order width does not yield fixed-parameter tractability for any of our problems on primitive task networks (Corollaries 21 and 22), at least under well-established complexity assumptions. This means that different structural parameterizations are required to obtain fixed-parameter tractability.
Further, Theorems 5 and 2 can already be seen to rule out fixed-parameter tractability for the considered problems on primitive task networks w.r.t. classical graph-theoretic measures such as treewidth Robertson and Seymour (1986), treedepth Nesetril and de Mendez (2012), and even the feedback edge number Uhlmann and Weller (2013); Ganian and Korchemna (2021). However, fixed-parameter tractability can still be achieved under the right parameterization: in Theorems 23 and 24, we obtain highly non-trivial fixed-parameter algorithms for Plan Validation, Plan Existence, and State Reachability when parameterized by the vertex cover number of the input primitive task network. Finally, for compound HTNs, we lift our general algorithmic meta-theorem for polynomial-time solvability (Theorem 16) to the setting of fixed-parameter tractability when additionally parameterizing by the “breadth” of an HTN, i.e., the maximum number of pairwise non-isomorphic networks that a task can be decomposed into (Theorem 25).
Related Work.
HTN planning has been extensively studied in the artificial intelligence research community and has found applications in areas such as healthcare García-Remesal (2008); González-Ferrer et al. (2011, 2013b); Bihlmaier et al. (2014), business processes González-Ferrer et al. (2013a); Ko et al. (2009, 2012), human-robot interaction Hayes and Scassellati (2016); Tewari and Persiani (2021), real-time games Ontañón and Buro (2015); Namiki et al. (2021); Soemers and Winands (2016), emergency response Liu et al. (2016); Zhao et al. (2017), and visualization Padia et al. (2019). For an overview of the current state of hierarchical planning at large, we refer the reader to the survey of Bercher, Alford, and Hoeller (2019) and the overview article of Georgievski and Aiello (2015).
The works most closely related to ours are those of Lin and Bercher (2023) and Olz, Biundo, and Bercher (2021), which also established some of the formalisms that we build on. In a similar vein, questions on the complexity and even computability of many tasks associated with HTNs have been treated in the literature Geier and Bercher (2011); Alford et al. (2015), where precursors of the abovementioned formal models were defined. Beyond these recent works, there are also classical investigations of computational questions related to HTN planning in slightly different settings than the ones we look at here, such as the early influential works by Erol, Hendler, and Nau (1994; 1996).
Albeit carried out in a different context, the work of Kronegger, Pfandler, and Pichler (2013) on propositional strips with negation (PSN) studies a problem ($k$-PSN) which is related to State Reachability. The main difference is that while PSN includes a broader definition of actions, it does not consider any primitive or hierarchical network structure whatsoever. Finally, we note that the computational complexity of planning has also been studied in a variety of other settings beyond HTNs Bylander (1994); Bäckström et al. (2012, 2013); Bäckström et al. (2015).
2 Preliminaries
For a positive integer $n$, let $[n]={\{1,\ldots,n\}}$. We employ standard graph-theoretic terminology Diestel (2012). The underlying undirected graph of a digraph $D$ is the graph obtained by replacing each arc in $D$ with an edge (discarding loops and multiple edges).
For some $V^{\prime}\subseteq V$, the subgraph induced by $V^{\prime}$ is $G[V^{\prime}]=(V^{\prime},E\cap(V^{\prime}\times V^{\prime}))$. A walk in $G$ is a sequence of vertices $v_{1},\ldots,v_{\ell}$ such that $(v_{i},v_{i+1})\in E$ for all $i\in[\ell-1]$. As the walk contains $\ell-1$ edges, we say that it has length $\ell-1$. A walk is a cycle if $v_{1}=v_{\ell}$. A graph is acyclic if it does not contain any cycles. A (simple) path is a walk such that $v_{i}\neq v_{j}$ for all $i\neq j,i,j\in[\ell]$. A simple cycle is a walk where this restrictions holds except for $v_{1}$ and $v_{\ell}$.
A multigraph is a graph that allows for multiple edges between a pair of vertices. Formally, we define a multigraph by $G=(V,E,\beta\colon E\rightarrow(V\times V))$. A vertex walk in $G$ is a sequence of vertices $v_{1},\ldots,v_{\ell}$ such that there exists $e_{i}\in E$ with $\beta(e_{i})=(v_{i},v_{i+1})$ for all $i\in[\ell-1]$. An edge walk in $G$ is a sequence of edges $e_{1},\ldots,e_{\ell}$ such that $e_{i}\in E$ for all $i\in[\ell]$ and if $\beta(e_{i})=(v,v^{\prime})$, then $\beta(e_{i+1})=(v^{\prime},v^{\prime\prime})$ for some $v^{\prime\prime}\in V$. Note that an edge walk corresponds to a vertex walk by considering the vertices it passes through. We transfer the notion of (simple) vertex paths and cycles and (simple) edge paths and cycles from non-multigraphs, where an edge walk is simple if and only if its corresponding vertex walk is simple.
The partial order width $w$ of a partially ordered set $\lesssim$ is the maximum number of pairwise-disjoint maximal chains contained in $\lesssim$ or, equivalently, the size of its largest antichain Dilworth (1950). An element in $\lesssim$ is isolated if it is incomparable to every other element in the set. We define the generalized partial order width of a partially ordered set $\lesssim$ as the maximum number of pairwise-disjoint maximal chains contained in $\lesssim$ which are not isolated elements (or, equivalently, the size of the largest antichain not containing any isolated element).
Hierarchical Task Networks.
Our formalization of HTNs follows the terminology employed in recent works Olz et al. (2021); Lin and Bercher (2023). The domain of a planning problem is a tuple $(F,A,{\cal C},\delta,M)$, where $F$ is a finite set of propositions, $A$ is a set of actions (or primitive task names), and ${\cal C}$ is a set of compound task names such that $A\cap{\cal C}=\emptyset$. The function $\delta\colon A\rightarrow 2^{F}\times 2^{F}\times 2^{F}$ maps each action to a set $\operatorname{prec}(a)$ of preconditions, a set $\operatorname{del}(a)$ of propositions to be removed, and a set $\operatorname{add}(a)$ of propositions to be added to the current state. Lastly, $M$ represents the decomposition methods and is a mapping from each $c\in{\cal C}$ to a set $Z$ of task networks (defined below) which $c$ can be decomposed into.
A task network in the domain $(F,A,{\cal C},\delta,M)$ is defined as a tuple $(T,\prec^{+},\alpha)$ that consists of a set $T$ of task identifiers with a partial order $\prec^{+}\ \subseteq\colon T\times T$ and a function $\alpha\colon T\rightarrow A\cup{\cal C}$. The usual representation considered for a task network is its cover graph $D_{\prec}=(T,\prec)$, where $\prec$ is the cover of $\prec^{+}$ (in particular, $(a,b)\in\prec$ if and only if $a\prec^{+}b$ and there is no element $c$ such that $a\prec^{+}c\prec^{+}b$). When analyzing the structure of directed graphs, it is often useful to also consider their underlying undirected graph Ordyniak and Szeider (2013); Ganian et al. (2021); we hence use $G_{\prec}$ to denote the underlying undirected graph of $D_{\prec}$. An example of a simple task network is provided in Fig. 1.
Action $\operatorname{prec}$ $\operatorname{del}$ $\operatorname{add}$
$a_{1}$emptyset $\emptyset$ ${\{1\}}$
$a_{2}$emptyset $\emptyset$ ${\{2\}}$
$a_{3}$${\{2\}}$${\{1\}}$emptyset
Figure 1: The digraph $D_{\prec}$ for a primitive HTN with tasks $T={\{t_{1},t_{2},t_{3},t_{4}\}}$, actions $A={\{a_{1},a_{2},a_{3}\}}$, propositions $F=\{1,2\}$, and initial state $s_{0}=\emptyset$. Each task’s action is given in its node. A directed edge from a task $t$ to a task $t^{\prime}$ corresponds to $(t,t^{\prime})\in\prec^{+}$.
A set of tasks $T^{\prime}\subseteq T$ is a chain if $\prec^{+}$ defines a total order on $T^{\prime}$, and an antichain if all tasks in $T^{\prime}$ are pairwise incomparable w.r.t. $\prec^{+}$. Two task networks $tn=(T,\prec^{+},\alpha)$ and $tn^{\prime}=(T^{\prime},{\prec^{+}}^{\prime},\alpha^{\prime})$ are isomorphic if there is a bijection $\phi\colon T\rightarrow T^{\prime}$ such that $(t,t^{\prime})\in\prec^{+}$ if and only if $(\phi(t),\phi(t^{\prime}))\in{\prec^{+}}^{\prime}$ and $\alpha(t)=\alpha^{\prime}(\phi(t))$ for all $t,t^{\prime}\in T$. Intuitively, two task networks are isomorphic if they describe the same structure, but using different task names.
A task $t$ is primitive if $\alpha(t)\in A$ and compound otherwise. A task network is primitive if it contains only primitive tasks. Non-primitive task networks can be decomposed into primitive ones by applying the decomposition methods in $M$. More precisely, let there be a task network $tn=(T,\prec^{+},\alpha)$, $t\in T$, and $m=(\alpha(t),tn_{m})\in M$. Let $tn_{m}^{\prime}=(T_{m}^{\prime},{\prec_{m}^{+}}^{\prime},\alpha_{m}^{\prime})$ with $T_{m}^{\prime}\cap T=\emptyset$ being isomorphic to $tn_{m}$. Then, $t$ in $tn$ can be decomposed by $m$ to obtain the new task network $tn^{\prime}=(T^{\prime},{\prec^{+}}^{\prime},\alpha^{\prime})$ with
$T^{\prime}$ $=T\setminus{\{t\}}\cup T_{m}^{\prime},$
${\prec^{+}}^{\prime}$ $=\prec^{+}\cap(T^{\prime}\times T^{\prime})\cup{\prec^{+}_{m}}^{\prime}\cup\{(t_{1},t_{2})\mid(t_{1},t)\in\prec^{+},$
$t_{2}\in T_{m}^{\prime}\}\cup{\{(t_{2},t_{1})\mid(t,t_{1})\in\prec^{+},t_{2}\in T_{m}^{\prime}\}},$
$\alpha^{\prime}$ $=\alpha\setminus{\{(t,\alpha(t))\}}\cup\alpha_{m}^{\prime}.$
Intuitively, decomposing a task $t$ to a subnetwork $tn_{m}$ means replacing $t$ by $tn_{m}$ such that each task in $tn_{m}$ is connected to tasks outside $tn_{m}$ in the same way $t$ was before. The extra step using the isomorphic $tn_{m}^{\prime}$ is required to ensure unique task identifiers after the decomposition. An example is provided in Fig. 2.
Figure 2: Example for the decomposition of a compound task network. Let $tn$ be the compound task network that is depicted in Fig. 1, where $t_{3}$ is a compound task instead of a primitive task with action $a_{1}$. If $(\alpha(t_{3}),tn_{m})\in M$, then $tn$ decomposes to $tn^{\prime}$.
An HTN planning problem is a problem whose input includes a domain $(F,A,{\cal C},\delta,M)$, an initial task network $tn$ from that domain, and an initial state $s_{0}\subseteq F$. An action $a$ is executable in a state $s\subseteq F$ if $\operatorname{prec}(a)\subseteq s$. If executed, it changes the state $s$ to $s^{\prime}=(s\setminus\operatorname{del}(a))\cup\operatorname{add}(a)$. A plan is a sequence of actions. A plan is executable from an initial state $s_{0}$ if all actions in the plan can be executed in sequence (one after the other) starting from $s_{0}$. A sequence $t_{1},t_{2},\ldots,t_{\ell}$ of tasks is executable in a task network $tn$ if $tn$ can be decomposed into a primitive task network $tn^{\prime}$ such that the plan $\alpha(t_{1}),\ldots,\alpha(t_{\ell})$ is executable, $t_{i}\neq t_{j}$, and $t_{i}\not\prec^{+}t_{j}$ if $i>j$. For a sequence $p$ of tasks or actions, we refer to its $i^{\text{th}}$ element by $p[i]$ and to its length by $|p|$.
Example 1.
Consider the primitive HTN depicted in Fig. 1. The task sequence $t_{1},t_{2},t_{4}$ is not executable since $(t_{3},t_{4})\in\prec^{+}$ means that $t_{3}$ must be executed before $t_{4}$. The task sequence $t_{1},t_{3}$, $t_{4}$ is not executable either, as after executing $t_{1}$ and $t_{3}$, the state is ${\{1\}}$, so the precondition of $a_{3}$ is not fulfilled when executing $t_{4}$. On the other hand, the task sequence $t_{1},t_{2},t_{3},t_{4}$ is executable. It conforms with the partial order, and for the corresponding action sequence $a_{1},a_{2},a_{1},a_{3}$, after executing any prefix of the sequence, the state is a superset of the precondition of the next action.
We are now ready to define our problems of interest.
Plan Validation
Input: A task network $tn=(T,\prec^{+},\alpha)$ over some domain $(F,A,{\cal C},\delta,M)$, an initial state $s_{0}\subseteq F$, and a plan $p$ over $A$.
Question: Is there a task sequence that corresponds to plan $p$, is executable in $tn$ from $s_{0}$, and executes all the tasks in $tn$?
As mentioned in the introduction, authors have also sometimes considered a version of Plan Validation where we are provided a sequence of tasks on the input as opposed to a plan. We remark that this would give rise to a problem which is trivially linear-time solvable on all primitive task networks, and is also fully captured by our algorithmic meta-theorems (Theorems 16 and 25) on compound networks. In view of this, we exclude this problem from our further considerations.
Plan Existence
Input: A task network $tn$ over some domain $(F,A,{\cal C},\delta,M)$ and an initial state $s_{0}\subseteq F$.
Question: Is there an executable plan starting in $s_{0}$ that executes all the tasks in $tn$?
Action Executability
Input: A task network $tn$ over some domain $(F,A,{\cal C},\delta,M)$, an initial state $s_{0}\subseteq F$, and a multiset $S$ of actions in $A$.
Question: Is there a plan (not necessarily using all available tasks) that is executable in $tn$ from $s_{0}$ and that covers $S$?
Action Executability is a generalization of Plan Existence in the sense that, on primitive task networks, Plan Existence simply equals Action Executability if $S$ is the multiset of actions obtained by mapping each task by $\alpha$. On general networks, any Plan Existence instance can be reduced to Action Executability by adding a new task $t$ with new action $a$, letting $S={\{a\}}$, and adding $(t^{\prime},t)$ to $\prec$ for each sink $t^{\prime}$ in the task network. As this may change some structural parameters associated to the task network, alternatively one can create a new task $t_{t^{\prime}}$ with action $a$ for each sink $t^{\prime}$, only add the edges $(t^{\prime},t_{t^{\prime}})$ to $\prec$, and let $S$ contain $a$ as many times as there are sinks. Therefore, algorithmic and hardness results transfer in the respective directions; in our setting, we establish all of our hardness results for the “simpler” Plan Existence, while all of our algorithms can deal with the “harder” Action Executability.
State Reachability
Input: A task network $tn$ over some domain $(F,A,{\cal C},\delta,M)$, an initial state $s_{0}\subseteq F$, and a target state $s_{g}\subseteq F$.
Question: Is there a plan (not necessarily using all available tasks) that is executable in $tn$ from $s_{0}$ and that results in a state $s\supseteq s_{g}$?
We remark that, while Plan Validation, Plan Existence, Action Executability, and State Reachability are stated as decision problems for complexity-theoretic purposes, all of our obtained algorithms are constructive and can output a task sequence or plan as a witness. For each of these problems, we define the input size to simply be the sum of the sizes of elements occurring on the input; for example, for an instance ${\cal I}$ of Plan Validation this would be $|{\cal I}|=|T|+|\prec^{+}|+|\alpha|+|F|+|A|+|{\cal C}|+|\delta|+|M|$.
For some of our algorithms, we employ a state transition graph. For a set of actions $A$ and an initial state $s_{0}$, we define the state transition graph to be a multigraph ${\cal G}=({\cal S},{\cal T},\beta)$ that describes the set ${\cal S}$ of states reachable from $s_{0}$ and how actions transform one state into another. We construct ${\cal G}$ inductively as follows. First, initialize ${\cal S}={\{s_{0}\}}$ and ${\cal T}=\emptyset$. Then, update from $s_{0}$. To update from a state $s$, iterate through all actions $a\in A$. If $\operatorname{prec}(a)\subseteq s$, let $s^{\prime}=(s\setminus\operatorname{del}(a))\cup\operatorname{add}(a)$. If $s^{\prime}\notin{\cal S}$, add $s^{\prime}$ to ${\cal S}$ and update from $s^{\prime}$. In any case, add $e=(a,s,s^{\prime})$ to ${\cal T}$ and let $\beta(e)=(s,s^{\prime})$. The computation is completed once no more updates occur. An example is provided in Fig. 3. This construction takes at most $\mathcal{O}\!\left(|{\cal I}|^{2}\right)$ time. Observe that ${\cal G}$ contains exactly all states that $s_{0}$ can be transformed into by applying sequences of actions in $A$. Further, a pair of states is connected by directed edges representing all actions that transform one of the states into the other. It is known that the number of states $|{\cal S}|$ can be upper-bounded by $|A|!\cdot 2^{|A|}$ Kronegger et al. (2013).
Action Prop. set $\operatorname{prec}$ $\operatorname{del}$ $\operatorname{add}$
$a_{1}$emptyset $\emptyset$ ${\{1\}}$
$a_{2}$${\{1\}}$emptyset ${\{2\}}$
$a_{3}$${\{2\}}$${\{2,3\}}$emptyset
$a_{4}$${\{1\}}$emptyset ${\{2,3\}}$
$a_{5}$${\{2\}}$emptyset ${\{3\}}$
Figure 3: Example of a state transition graph with initial state $s_{0}=\emptyset$ and actions $A={\{a_{1},a_{2},a_{3},a_{4},a_{5}\}}$. An edge labeled with $a_{i}$ that connects some state $s$ to some state $s^{\prime}$ refers to the edge $(a_{i},s,s^{\prime})$ in the state transition graph. An edge labeled with multiple actions refers to multiple edges, one for each action, in the multigraph.
Parameterized Complexity.
Parameterized complexity Downey and Fellows (2013); Cygan et al. (2015) analyzes the running time of algorithms with respect to a parameter $k\in\mathbb{N}$ and input size $n$. The high-level idea is to find a parameter that describes the structure of the instance such that the combinatorial explosion can be confined to this parameter. In this respect, the most favorable complexity class is FPT (fixed-parameter tractable) which contains all problems that can be decided by an algorithm running in $f(k)\cdot n^{\mathcal{O}(1)}$ time, where $f$ is a computable function. Algorithms with this running-time are called fixed-parameter algorithms.
Showing that a problem is hard for the complexity classes ${\textsf{W}}[1]$ or ${\textsf{W}}[2]$ rules out the existence of a fixed-parameter algorithm under the well-established assumption that ${\textsf{W}}[1]\neq{\textsf{FPT}}$. This is done via a parameterized reduction Downey and Fellows (2013); Cygan et al. (2015) from some known ${\textsf{W}}[1]$-hard (${\textsf{W}}[2]$-hard, respectively) problem. A parameterized reduction from a parameterized problem $\mathcal{P}$ to a parameterized problem $\mathcal{Q}$ is a function:
  • which maps yes-instances to yes-instances and no-instances to no-instances,
  • which can be computed in $f(k)\cdot n^{\mathcal{O}(1)}$ time, where $f$ is a computable function, and
  • where the parameter of the output instance can be upper-bounded by some function of the parameter of the input instance.
3 Setting the Stage: The State Space
As a basic precondition for solving our problems on more advanced task networks, we need to ensure tractability at least on “trivial” primitive task networks where $\prec^{+}=\emptyset$. Unfortunately, in their most general form, all but the first of our problems of interest turn out to be intractable even on these degenerate networks via a simple reduction from 3-SAT.
Theorem 1.
State Reachability and Plan Existence are NP-hard on primitive task networks with $\prec^{+}\ =\emptyset$.
Proof.
We give a reduction from 3-SAT to State Reachability. Given an instance $\phi$ of 3-SAT, we create an instance $(tn,s_{0},s_{g})$ of State Reachability as follows. For each clause $C_{j}$ in $\phi$, $j\in[m]$, we create a proposition $C_{j}$ in $F$. For each variable $x_{i}$ in $\phi$, $i\in[n]$, we create a proposition $x_{i}$ in $F$, as well as two actions $a_{x_{i}}$, $a_{\overline{x}_{i}}$ and two tasks $t_{x_{i}}$, $t_{\overline{x}_{i}}$ that are mapped to $a_{x_{i}}$ and $a_{\overline{x}_{i}}$, respectively. For all $i\in[n]$, let $C_{x_{i}}$ ($C_{\overline{x}_{i}}$, respectively) be the set of all propositions of the form $C_{j}$ that correspond to clauses in $\phi$ that are satisfied by setting the variable $x_{i}$ to TRUE (FALSE, respectively). For all $i\in[n]$, we set $\operatorname{prec}(a_{x_{i}})=\operatorname{prec}(a_{\overline{x}_{i}})=\operatorname{del}(a_{x_{i}})=\operatorname{del}(a_{\overline{x}_{i}})=\{x_{i}\}$, $\operatorname{add}(a_{x_{i}})=C_{x_{i}}$, and $\operatorname{add}(a_{\overline{x}_{i}})=C_{\overline{x}_{i}}$. Finally, we set $s_{0}=\{x_{1},\ldots,x_{n}\}$ and $s_{g}=\{C_{1},\ldots,C_{m}\}$. Note that $tn$ is primitive with $\prec^{+}=\emptyset$. The equivalence between the two instances is easy to see. Indeed, since $s_{0}=\{x_{1},\ldots,x_{n}\}$ and $\operatorname{prec}(a_{x_{i}})=\operatorname{prec}(a_{\overline{x}_{i}})=\operatorname{del}(a_{x_{i}})=\operatorname{del}(a_{\overline{x}_{i}})=\{x_{i}\}$, only one of the two tasks $t_{x_{i}}$, $t_{\overline{x}_{i}}$ may be executed for each $i\in[n]$. Further, the execution of $t_{x_{i}}$ ($t_{\overline{x}_{i}}$, respectively) causes the execution of $a_{x_{i}}$ ($a_{\overline{x}_{i}}$, respectively), which adds, to the current state, all the propositions corresponding to clauses in $\phi$ that are satisfied by setting the variable $x_{i}$ to TRUE (FALSE, respectively). Finally, $s_{g}$ contains all the clause propositions. Hence, $\phi$ is satisfiable if and only if $(tn,s_{0},s_{g})$ is a yes-instance of State Reachability. We can also show that the same holds for Action Executability, even if $|S|=1$; indeed, it suffices to use the same construction with the addition of a new task $t_{a^{*}}$ that is mapped to a new action $a^{*}$ with $S=\{a^{*}\}$ and $\operatorname{prec}(a^{*})=\{C_{1},\ldots,C_{m}\}$. Finally, for Plan Existence, we begin by recalling that the problem is NP-hard on primitive task networks (Erol et al., 1996, Theorem 8). We can reduce from Plan Existence on primitive task networks to Plan Existence on primitive task networks with $\prec^{+}=\emptyset$ by encoding the partial order into the preconditions. Let there be a task network $tn=(T,\prec^{+},\alpha)$ with action set $A$ and set of propositions $F$. Then, construct a new task network $tn^{\prime}=(T,\emptyset,\alpha^{\prime})$ with action set $A^{\prime}$ and proposition set $F^{\prime}$ as follows. Let $F^{\prime}=F\cup{\{f_{t}\mid t\in T\}}$. For each $t\in T$, create an action $a_{t}\in A^{\prime}$ such that $\operatorname{prec}(a_{t})=\operatorname{prec}(\alpha(t))\cup{\{f_{t^{\prime}}\mid(t^{\prime},t)\in\prec^{+}\}}$, $\operatorname{del}(a_{t})=\operatorname{del}(\alpha(t))$, and $\operatorname{add}(a_{t})=\operatorname{add}(\alpha(t))\cup{\{f_{t}\}}$. Let $\alpha^{\prime}(t)=a_{t}$ for all $t\in T$. Observe that this way, a task in $tn^{\prime}$ can only be executed if the respective preconditions as defined for the original instance are fulfilled and all its predecessors in $\prec^{+}$ are executed. Thus, any task sequence is executable in $tn^{\prime}$ from some initial state $s_{0}$ if and only if it is executable in $tn$ from $s_{0}$. ∎
A crucial requirement for the reductions underlying Theorem 1 is that the state space (i.e., the number of states) can be large, and in particular not bounded by any constant. As we will see in the next sections, tractability can be achieved if the state space is assumed to be fixed (that is, bounded by a constant that is independent of the input). Hence, we will hereinafter analyze the affected problems (State Reachability, Plan Existence, and its generalization Action Executability) in the bounded state-space setting.
4 Primitive Task Networks
Our first set of results focuses on the classical complexity of the targeted problems on primitive task networks. We begin with a series of lower bounds which rule out polynomial-time algorithms even for some very simple “tree-like” networks.
The Intractability of Tree-like HTNs.
First, note that the intractability of Plan Validation on tree-like networks was already established by previous work of Lin and Bercher:
Theorem 2 (Lin and Bercher (2023), Theorem 3).
Plan Validation on primitive task networks is NP-hard even if $G_{\prec}$ is a forest of stars and $\operatorname{prec}(a)=\operatorname{del}(a)=\operatorname{add}(a)=\emptyset$ for all actions $a\in A$.
Further, Plan Validation remains intractable even if $\prec^{+}$ describes a set of total orders that are independent from one another (i.e., if $D_{\prec}$ is a disjoint union of paths). We establish this via a reduction from the Shuffle Product problem:
Shuffle Product
Input: Words $u,c_{1},\ldots,c_{w}$ over an alphabet $\Sigma$.
Question: Can $u$ be generated by interlacing the words in $c_{1},\ldots,c_{w}$, i.e., can the letters of $c_{1},\ldots,c_{w}$ be rearranged to form $u$ while preserving the order the letters appear in for each individual word $c_{i}$, $i\in[w]$?
Shuffle Product is known to be NP-hard even when restricted to the case where $|\Sigma|=2$ Warmuth and Haussler (1984).
Theorem 3.
Plan Validation is NP-hard, even when restricted to primitive task networks such that $\prec^{+}$ is a collection of chains, $|A|=2$, and $\operatorname{prec}(a)=\operatorname{del}(a)=\operatorname{add}(a)=\emptyset$ for all actions $a\in A$.
Proof.
We reduce from Shuffle Product with the alphabet $\Sigma={\{a,b\}}$. Let $A=\Sigma$. Construct a task for each letter in the words $c_{1},\ldots,c_{w}$. In particular, let the task $t_{i,j}$ refer to the $j^{\text{th}}$ letter of the word $c_{i}$, $i\in[w]$. Let $\prec^{+}$ be such that the tasks for each word form a chain in the respective order, i.e., let $\prec$ consist of the edges $(t_{i,j},t_{i,j+1})$, for all $i\in[w]$ and $j\in[|c_{i}|-1]$. Then, $\prec^{+}$ partitions $T$ into $w$ chains (see Fig. 4). Now, set $\alpha(t_{i,j})$ to the action representing the $j^{\text{th}}$ letter of $c_{i}$. Let the input plan $p$ be such that $p=u$. The constructed instance consists of $w$ chains and the construction takes polynomial time.
The correctness of the reduction follows from the fact that verifying whether the tasks in this Plan Validation instance can be arranged to match the input plan corresponds to interlacing the words to form $u$. ∎
A different reduction from Shuffle Product yields similar hardness results for the other problems; it will be useful to recall that hardness for Plan Existence directly carries over to hardness for Action Executability.
Theorem 4.
State Reachability and Plan Existence are NP-hard, even when restricted to primitive task networks such that $\prec^{+}$ is a collection of chains, $|A|\leq 5$, and the state transition graph has at most $4$ states.
Proof.
We reduce from Shuffle Product with the alphabet $\Sigma={\{a,b\}}$. Construct a task for each letter in the words $u,c_{1},\ldots,c_{w}$. In particular, for each word $x\in{\{u,c_{1},c_{2},\ldots,c_{w}\}}$, let the task $t_{x,j}$ refer to its $j^{\text{th}}$ letter. Let $T_{c}$ be the set of tasks created for letters in $c_{1},\ldots,c_{w}$, and $T_{u}$ the set of tasks $t_{u,j}$, $j\in|u|$. Let $\prec^{+}$ be such that the tasks for each word form a chain in the respective order, i.e., let $\prec$ consist of the edges $(t_{x,j},t_{x,j+1})$ for each $x\in{\{u,c_{1},\ldots,c_{w}\}}$ and $j\in[|x|-1]$. Further, add a task $t_{g}$ and let $(t_{u,|u|},t_{g})\in\prec$. Then, $\prec^{+}$ partitions $T$ into $w+1$ chains (see Fig. 4).
For the sake of visualization and the corresponding naming of actions and propositional states, we consider the chains for $c_{1},\ldots,c_{w}$ to be on some left side and the chain for $u$ to be on some right side. Let $A={\{a^{L}_{a},a^{L}_{b},a^{R}_{a},a^{R}_{b},a_{g}\}}$ with the preconditions and effects of these actions described in Table 1.
Action Prop. set $\operatorname{prec}$ $\operatorname{del}$ $\operatorname{add}$
[…]
Table 1: Preconditions and effects of actions in the proof of Thm. 4.
If the $j^{\text{th}}$ letter of $u$ is an $a$, then we set $\alpha(t_{u,j})=a^{R}_{a}$, and otherwise, $\alpha(t_{u,j})=a^{R}_{b}$. Similarly, for each $i\in[w]$, if the $j^{\text{th}}$ letter in $c_{i}$ is an $a$, then we set $\alpha(t_{c_{i},j})=a^{L}_{a}$, and otherwise, $\alpha(t_{u,j})=a^{L}_{b}$. Further, let $\alpha(t_{g})=a_{g}$ and $s_{0}={\{\text{LEFT}\}}$. For State Reachability, we let $s_{g}={\{\text{GOAL}\}}$.
[…]
For the other direction, suppose there is a task sequence solving the State Reachability or Plan Existence instance. Consider the first part of the solution sequence up to the task $t_{g}$. All the tasks in $T_{u}$ are executed before $t_{g}$, and $\prec^{+}$ ensures that they are executed in the order of their corresponding letters in $u$. Note that, by the effects and preconditions of the actions, the sequence starts with a task in $T_{c}$ and alternates between tasks in $T_{c}$ and $T_{u}$. Further, a task in $T_{c}$ is always followed by a task in $T_{u}$ that corresponds to the same letter. Thus, listing all the letters in the words $c_{1},\ldots,c_{w}$ in the order their corresponding tasks are executed in the task sequence, gives $u$. As $\prec^{+}$ ensures that the letters from each word $c_{i}$, $i\in[w]$, are listed in the order they appear in $c_{i}$, it is a yes-instance of Shuffle Product.
[…]
Figure 4: Constructed task network for reductions from Shuffle Product with words $c_{1},\ldots,c_{w}$ and $u$. Each word is represented by a chain of tasks where each task represents a letter in that word. The colors orange and blue represent the two different letters in the alphabet. For the reduction to Plan Validation (Theorem 3), the task network only consists of vertices on the LEFT side. For the reductions to State Reachability and Plan Existence (Theorem 4), a task’s action is defined by its color and whether it is on the LEFT or RIGHT side.
Having established the intractability of all three problems on collections of paths, it is natural to wonder whether State Reachability and Plan Existence can at least be solved on collections of stars. Here, too, we provide a negative answer via a reduction from the strongly NP-hard 3-Partition problem Garey and Johnson (1979):
3-Partition
Input: $m,B\in\mathbb{N}$ and $3m$ positive integers $n_{1},\ldots,n_{3m}$ such that $\sum_{i=1}^{3m}n_{i}=mB$.
Question: Can the integers $n_{i}$ be partitioned into $m$ triples such that each triple sums up to $B$?
Theorem 5.
State Reachability and Plan Existence are NP-hard on primitive task networks that are forests of stars, even if $|A|=8$ and the state transition graph has at most $9$ states.
Proof.
We reduce from 3-Partition. We employ the set of actions $A={\{a_{\emptyset},a_{\beta},a_{\hat{\beta}},a_{\gamma},a_{\hat{\gamma}},a_{\delta},a_{\hat{\delta}},a_{g}\}}$. The preconditions and effects of these actions are described in Figure 5, which also illustrates the construction of the following task network. For each $i\in[3m]$, create a star representing the integer $n_{i}$ as follows. Create a task $t_{i}$ and two task sets $T_{i},T_{i}^{\prime}$, each of cardinality $n_{i}$. Let $\alpha(t_{i})=a_{\emptyset}$. For all $t\in T_{i}$, let $\alpha(t)=a_{\hat{\beta}}$ and $(t,t_{i})\in\prec$. For all $t\in T_{i}^{\prime}$, let $\alpha(t)=a_{\gamma}$ and $(t_{i},t)\in\prec$. Next, create $m$ more gadgets representing the $m$ triples to be built, where each triple is represented by a star as follows. For all $i\in[m-1]$, the $i^{\text{th}}$ triple has a center vertex $t_{\delta,i}$, a set $T_{\hat{\gamma}}$ of $B$ tasks, and a set $T_{\beta}$ of $B$ tasks. Let $\alpha(t_{\delta,i})=a_{\delta}$. For all $t\in T_{\hat{\gamma}}$, let $\alpha(t)=a_{\hat{\gamma}}$ and $(t,t_{\delta,i})\in\prec$. For all $t\in T_{\beta}$, let $\alpha(t)=a_{\beta}$ and $(t_{\delta,i},t)\in\prec$. The star representing the last triple differs from the above construction. The last triple has a task $t_{\delta,m}$ and a set $T_{\hat{\gamma}}$ of $B$ tasks as described above, but no set $T_{\beta}$. To initiate things, there is also an independent set of $B$ tasks all with action $a_{\beta}$. Lastly, create a star to count the completed triples. This star consists of a center vertex $t_{g}$ and a set $T_{\hat{\delta}}$ of $m$ leaf-tasks. Let $\alpha(t_{g})=a_{g}$ and, for each $t\in T_{\hat{\delta}}$, let $\alpha(t)=a_{\hat{\delta}}$ and $(t,t_{g})\in\prec$. Let the initial state be $s_{0}={\{\hat{\beta},\hat{\gamma}\}}$. For State Reachability, let $s_{g}={\{\text{GOAL}\}}$. We prove that these instances are yes-instances if and only if the 3-Partition instance is a yes-instance.
Action Prop. set $\operatorname{prec}$ $\operatorname{del}$ $\operatorname{add}$
$a_{\emptyset}$emptyset $\emptyset$ $\emptyset$
$a_{\beta}$${\{\hat{\beta}\}}$${\{\hat{\beta}\}}$${\{\beta\}}$
$a_{\hat{\beta}}$${\{\beta\}}$${\{\beta\}}$${\{\hat{\beta}\}}$
$a_{\gamma}$${\{\hat{\gamma}\}}$${\{\hat{\gamma}\}}$${\{\gamma\}}$
$a_{\hat{\gamma}}$${\{\gamma\}}$${\{\gamma\}}$${\{\hat{\gamma}\}}$
$a_{\delta}$emptyset $\emptyset$ ${\{\delta\}}$
$a_{\hat{\delta}}$${\{\delta\}}$${\{\delta\}}$emptyset
[…]
Figure 5: Actions and constructed task network for the reduction in the proof of Theorem 5. The label at each task $t$ either refers to both its name and action ($t:\alpha(t)$) or, for legibility, just its action.
Suppose there is a 3-Partition of the integers $n_{i}$. We build a task sequence that enumerates these triples. Let $(n_{x},n_{y},n_{z})$ be the first triple in the solution. Initially, there are $B$ tasks with action $a_{\beta}$ that are executable. Alternate executing these tasks and leaf-tasks with action $a_{\hat{\beta}}$ of the stars with centers $t_{x}$, $t_{y}$, and $t_{z}$. This is possible as $n_{x}+n_{y}+n_{z}=B$ and the preconditions and effects of $a_{\beta}$ and $a_{\hat{\beta}}$ enforce this alternating behavior. Now, execute $t_{x}$, $t_{y}$, and $t_{z}$. Then, alternate between executing one of the $B$ tasks with action $a_{\gamma}$ in the stars of $t_{x}$, $t_{y}$, and $t_{z}$, and the $B$ tasks with action $a_{\hat{\gamma}}$ in the star of $t_{\delta,1}$. Then, execute $t_{\delta,1}$, followed by one of the not yet executed tasks of action $t_{\hat{\delta}}$. After executing this sequence, we are in a similar state as in the beginning. Once more, there are $B$ tasks with action $a_{\beta}$ executable. However, one of the tasks with action $a_{\hat{\delta}}$ and all of the tasks in the star with center $t_{\delta,1}$, as well as the three stars corresponding to $n_{x}$, $n_{y}$, and $n_{z}$ are executed. After repeating the above procedure for all $m$ triples, the only task remaining is $t_{g}$ (since the last star representing a triple does not have a set $T_{\beta}$) and it is executable, which results in a yes-instance for State Reachability and Plan Existence.
For the other direction, suppose we have a yes-instance of State Reachability or Plan Existence. This implies that there is a sequence executing $t_{g}$. This task is only executable after the execution of all $m$ tasks with action $a_{\hat{\delta}}$, which in turn requires the execution of all $m$ tasks $t_{\delta,i}$ for $i\in[m]$. Without loss of generality, let these tasks be executed in the order $t_{\delta,1},\ldots,t_{\delta,m}$. We define multisets $S_{i}$ for $i\in[m]$ as follows. Let $S_{1}={\{n_{j}\mid t_{j}\text{ was executed before }t_{\delta,1}\}}$ and, for $i\in[m-1]$, let $S_{i+1}={\{n_{j}\mid t_{j}\text{ was executed after $t_{\delta,i}$ and before }t_{\delta,i+1}\}}$. Observe that $P={\{S_{i}\mid i\in[m]\}}$ partitions the integers $n_{1},\ldots,n_{3m}$. In order to execute a task $t_{j}$ for $j\in[3m]$, first its $n_{j}$ in-neighbors and, by their preconditions, $n_{j}$ tasks of action $a_{\beta}$ have to be executed. The execution of the task $t_{j}$ then enables the execution of $n_{j}$ tasks of action $a_{\gamma}$. Note that, for the execution of $t_{\delta,1}$, exactly $B$ tasks with action $a_{\beta}$ are available, and hence, $\sum_{n\in S_{1}}\leq B$, and the execution of $B$ tasks with action $a_{\gamma}$ is required, and hence, $\sum_{n\in S_{1}}\geq B$. Thus, $\sum_{n\in S_{1}}=B$ and, after the execution of $t_{\delta,1}$, there are once more exactly $B$ tasks of action $a_{\beta}$ available and each of the not yet executed tasks $t_{j}$ still has $n_{j}$ not yet executed in-neighbors. Hence, we can repeatedly apply our argument and obtain $\sum_{n\in S_{i}}=B$ for all $i\in[m]$. Thereby, $P$ is a 3-Partition. Lastly, observe that, for each reachable state $s$ for which $\text{GOAL}\notin s$, either $\beta\in s$ or $\hat{\beta}\in s$, either $\gamma\in s$ or $\hat{\gamma}\in s$, and either $\delta\in s$ or $\delta\notin s$. Since $s={\{\text{GOAL}\}}$ when $\text{GOAL}\in s$, this gives $k\leq 2^{3}+1=9$ reachable states. ∎
Chains and Antichains.
Given the seemingly discouraging lower bounds obtained so far in this section, one might rightfully wonder whether there are reasonable structural restrictions on the network which can be exploited to obtain polynomial-time algorithms. In this subsection, we chart the route towards identifying these by first establishing the tractability of all considered problems (under the state-space restrictions justified in Section 3) in the extremes of $\prec^{+}$, that is, where $T$ is a chain (i.e., total order) or an antichain ($\prec^{+}\ =\emptyset$).
We begin by showing that the problems are easily solved on chains. Recall from our discussion in Section 2 that on primitive task networks, polynomial-time solvability results for Action Executability carry over to Plan Existence.
Observation 6.
Plan Validation, State Reachability, and Action Executability are linear-time solvable on primitive task networks such that $\prec^{+}$ defines a total order.
Proof.
Note that $\prec^{+}$ defines the exact order in which the tasks in $T$ must be executed and that in order to execute a certain task, all tasks that precede it in the order must be executed first. It thus suffices to iteratively execute the tasks in the order given by $\prec^{+}$ and at each step check whether the respective action given by $\alpha$ can be executed in the current state and otherwise abort. For State Reachability and Action Executability, we have a yes-instance if after some step the target state is reached or the action set is covered. For Plan Validation with plan $p$, we have a yes-instance if, for the $i^{\text{th}}$ task $t_{i}$, we have $\alpha(t_{i})=p[i]$ for all $i\in[|T|]$. ∎
It is also trivial to solve Plan Validation on antichains.
Observation 7.
Plan Validation is linear-time solvable on primitive task networks such that $\prec^{+}=\emptyset$.
Proof.
Since $\prec^{+}=\prec=\emptyset$, there are no precedence constraints on the order the tasks may be executed with respect to task names, and thus, we can simply map the tasks to their corresponding multiset of actions. Deciding whether a plan is executable in the task network is achieved by first checking whether the number of action occurrences in the sequence matches the multiset of actions and then simulating the sequence to check whether the preconditions of each action are fulfilled. ∎
Handling antichains for the remaining problems already requires more effort. For State Reachability, we can obtain an algorithm via an analysis of how a hypothetical solution traverses the state graph.
Theorem 8.
State Reachability is polynomial-time solvable on primitive antichain task networks with constantly many states.
Proof.
Let $k$ be the number of states, $s_{0}$ be the initial state and $s_{g}$ the target state. We begin by computing the state transition graph ${\cal G}$ of ${\cal I}$ in at most quadratic time. In ${\cal G}$, if there are more than $k$ edges between any pair of states, arbitrarily select $k$ of them and discard the rest. Then, consider all simple edge paths from $s_{0}$ to $s_{g}$. Note that there are $\mathcal{O}\!\left(2^{k}\cdot k^{k-1}\right)$ simple edge paths, as there are $\mathcal{O}\!\left(2^{k}\right)$ simple vertex paths, each of length at most $k-1$, and each step might choose between up to $k$ different edges. For each simple edge path, check whether there is any action that is used more times in the path than there are tasks of that action in $T$. If there exists a simple edge path for which the number of required tasks does not exceed the number of available ones, then give a positive answer, and otherwise return that it is a no-instance.
If such a path is found, the instance is indeed a yes-instance as there is a task sequence that results in a path to $s_{g}$ in ${\cal G}$. For the other direction, let there be a task sequence that leads to the state $s_{g}$. Consider the subsequence obtained by exhaustively removing cycles in ${\cal G}$ from that sequence. The resulting task sequence describes a simple edge path from $s_{0}$ to $s_{g}$. It might employ edges that were discarded from ${\cal G}$ by the algorithm. However, in this case there are $k$ other edges connecting the respective states in ${\cal G}$. Thus, one can choose to replace deleted edges by existing ones in such a way that each replacement action occurs at most once in the path, which ensures that there are enough tasks to actually walk this path. Hence, there is a path from $s_{0}$ to $s_{g}$ that only uses edges that are present in ${\cal G}$ after the deletion and for which the number of required tasks does not exceed the number of available ones. As the algorithm will find this path, it will correctly give a positive output.
Reducing the number of edges in ${\cal G}$ takes $\mathcal{O}\!\left(|A|\cdot k\right)$ time, while the number of tasks for each action can be counted in $\mathcal{O}\!\left(|T|\right)$ time. For each path of length at most $k$, it takes $\mathcal{O}\!\left(k\right)$ time to check whether there are enough tasks for each action in the path. This gives a total runtime of $\mathcal{O}\!\left(|{\cal I}|^{2}+|A|k+|T|+2^{k}\cdot k^{k}\right)$, where the middle two terms are always dominated by either the first or the last term. Hence, any instance ${\cal I}$ of State Reachability in which the task network is primitive and $\prec^{+}=\emptyset$ can be solved in $\mathcal{O}\!\left(|{\cal I}|^{2}+2^{k}\cdot k^{k}\right)$ time, where $k$ is the number of states. ∎
Observe that when trying to adapt the above approach for Plan Existence or, more generally, for Action Executability, we might now have to visit a state multiple times as multiple tasks or actions might be executable in that state alone. The number of times a state is visited is bounded by the size of the action set $S$, however this insight only yields a polynomial-time algorithm if $k+|S|$ is fixed. In particular, we note that for the reduction from Plan Existence to Action Executability, it may possibly happen that $|S|=\Theta(|T|)$. Nevertheless, we obtain a polynomial-time algorithm for both problems on instances with fixed $k$ by a more complex approach which exploits the fact that sufficiently long walks in the state transition graph will revisit the same simple cycle multiple times.
Before we present the algorithms, we will need to introduce two reduction rules and the notion of equivalent actions.
Definition 4.1.
For an HTN planning instance with state transition graph ${\cal G}=({\cal S},{\cal T},\beta)$, let $a_{1},a_{2}\in A$ and, for all $s,s^{\prime}\in{\cal S}$, let $(a_{1},s,s^{\prime})\in{\cal T}$ if and only if $(a_{2},s,s^{\prime})\in{\cal T}$. Then, $a_{1}$ and $a_{2}$ are equivalent actions.
We note that there are at most $(k+1)^{k}$ equivalence classes because each class is identified by a vector $D$ of length $k=|{\cal S}|$ over ${\cal S}\cup{\{\emptyset\}}$. Thereby, $D[i]=s$ is to be interpreted as the equivalence class transforms the $i^{\text{th}}$ state into state $s$, and $D[i]=\emptyset$ implies the class does not transform the $i^{\text{th}}$ state into anything as its preconditions are not fulfilled. This allows us to define a reduction rule reducing the number of different actions. A reduction rule is safe if the rule preserves (or correctly identifies) yes- and no-instances.
For any primitive task network $tn=(T,\prec^{+},\alpha)$ and action $a\in A$, let $T_{a}={\{t\in T\mid\alpha(t)=a\}}$ and $\max(a)=|T_{a}|$. For Action Executability instances with action multiset $S$, let $\min(a)$ denote the number of occurrences of $a$ in $S$. Note that a primitive Action Executability instance has a solution if and only if there is a task sequence that is executable in $tn$ and $s_{0}$ in which each action $a$ occurs at least $\min(a)$ and at most $\max(a)$ times. We now provide the reduction rules we will use in our algorithm for Action Executability.
Reduction Rule R0:
For a primitive Action Executability instance, if there is an action $a\in A$ such that $\min(a)>\max(a)$, then output “no”.
Reduction Rule R1:
Consider a primitive Action Executability instance such that Reduction Rule R0 cannot be applied. Let $a_{1},a_{2}\in A$ with $a_{1}\neq a_{2}$ be equivalent actions. If, for all $t_{1}\in T_{a_{1}},t_{2}\in T_{a_{2}},t\in T$, the following holds:
  • $(t,t_{1})\in\prec^{+}$ if and only if $(t,t_{2})\in\prec^{+}$ and
  • $(t_{1},t)\in\prec^{+}$ if and only if $(t_{2},t)\in\prec^{+}$,
then change $\alpha$ such that, for all $t_{2}\in T_{a_{2}}$, now $\alpha(t_{2})=a_{1}$, and replace all occurrences of $a_{2}$ in $S$ by $a_{1}$.
Lemma 9.
Reduction Rule R1 is safe.
Proof.
Let $\max_{r}(a)$ and $\min_{r}(a)$ be defined as above, but for the reduced instance. Hence, $\max_{r}(a_{1})=\max(a_{1})+\max(a_{2})$ and $\min_{r}(a_{1})=\min(a_{1})+\min(a_{2})$. Thus, any task sequence that solves the original instance also solves the reduced instance. Now assume there is a task sequence that solves the reduced instance. If the same task sequence solves the original instance, we are done. Otherwise, let $x_{1}$ and $x_{2}$ be the number of occurrences of $a_{1}$- and $a_{2}$-tasks, respectively, in the sequence resolving the reduced instance according to the original mapping $\alpha$. Note that $\max_{r}(a_{1})=\max(a_{1})+\max(a_{2})\geq x_{1}+x_{2}\geq\min(a_{1})+\min(a_{2})$. Due to the equivalence in their actions and neighborhoods in $\prec^{+}$, $a_{1}$- and $a_{2}$-tasks can be used interchangeably in the sequence. Thus, any $a_{1}$-task can be replaced by an $a_{2}$-task, and vice versa, which, due to the above inequalities, makes it possible to find a valid solution for the original instance, i.e., one in which $\max(a_{1})\geq x_{1}\geq\min(a_{1})$ and $\max(a_{2})\geq x_{2}\geq\min(a_{2})$. ∎
Note that an instance of Action Executability with $|S|>|T|$ is a trivial no-instance, so we assume $|S|\leq|T|$. By combining the above reduction rules with layered branching techniques and integer programming, we prove:
Theorem 10.
Action Executability is polynomial-time solvable on primitive antichain task networks with constantly many states.
Proof.
We begin by exhaustively applying Reduction Rules R0 and R1, which takes at most $\mathcal{O}(|{\cal I}|^{3})$ time. Afterwards, we have either correctly solved the instance, or—since the network is an antichain—obtained a new equivalent instance where the number of actions is bounded by a function of the number $k$ of states, and in particular $|A|\leq(k+1)^{k}$.
The algorithm now branches over the at most $2^{k}\cdot|A|^{k}$ simple edge paths in ${\cal G}$ that start at $s_{0}$. Let $C$ be the set of distinct simple edge cycles in ${\cal G}$, where two cycles are considered distinct if and only if they do not contain the exact same vertices, i.e., cycles obtained from other cycles by inverting their directions are not counted as different cycles. Then, $|C|\leq 2^{k}\cdot|A|^{k}$. The algorithm branches on all possible subsets $C^{\prime}\subseteq C$ to decide which of these cycles are used at least once by a solution sequence. To ensure that there is a walk, set all cycles in $C^{\prime}$ as unmarked and exhaustively mark cycles that either share a vertex with the chosen path or an already marked cycle. Proceed with $C^{\prime}$ only if all cycles are marked by this procedure.
We write an Integer Linear Program (ILP) as follows. Let the ILP have a variable $x_{c}$ for every $c\in C^{\prime}$. For each action $a$ and $c\in C^{\prime}$, let $c(a)$ denote the number of occurrences of $a$ in $c$. Further, let $p(a)$ denote the number of occurrences of $a$ in the chosen path. For each $a\in A$, require the two constraints
$p(a)+c_{1}(a)x_{c_{1}}+c_{2}(a)x_{c_{2}}+\dots c_{|C^{\prime}|}(a)x_{c_{|C^{\prime}|}}$ $\leq\max(a)$
$p(a)+c_{1}(a)x_{c_{1}}+c_{2}(a)x_{c_{2}}+\dots c_{|C^{\prime}|}(a)x_{c_{|C^{\prime}|}}$ $\geq\min(a),$
where $c_{1},\ldots,c_{|C^{\prime}|}$ are the elements of $C^{\prime}$. Further, for all $c\in C^{\prime}$, add a constraint $x_{c}\geq 1$ to ensure that each of the selected cycles is walked at least once. The algorithm decides that the given instance is a yes-instance if and only if there is a solution to the ILP.
If the ILP has a solution, construct a plan $p$ as follows. Start with the chosen path. For each cycle $c\in C^{\prime}$, add $x_{c}$ repetitions of that cycle at the first time a vertex of $c$ appears in the path. As the cycles end in the same state they started in, the resulting action sequence is executable. Possibly, some cycles cannot be added directly, but only after other cycles have been added. Nevertheless, as all cycles in $C^{\prime}$ were marked by the algorithm, and the ILP ensures that each cycle in $C^{\prime}$ is executed at least once, all cycles can be added in this way. Further, the ILP ensures that each action $a$ appears at least $\min(a)$ and at most $\max(a)$ times in the resulting executable action sequence. Thus, the given instance is a yes-instance.
For the other direction, suppose there is an executable plan $p$ such that each action $a$ occurs at least $\min(a)$ and at most $\max(a)$ times in $p$. Consider an initially empty set of cycles $C^{*}$. Then, find a substring of $p$ that describes a simple edge cycle in ${\cal G}$. Add this cycle to $C^{*}$ while disregarding its starting vertex and its direction, remove the respective substring from $p$, and repeat until no more such cycles exist. Observe that $p$ then describes a simple path in ${\cal G}$, and $C^{*}$ exclusively contains simple edge cycles in ${\cal G}$. There is a branch of the algorithm that chooses this simple path and has $C^{\prime}=C^{*}$. Further, as the cycles were iteratively removed from a walk in the state transition graph, the algorithm will properly mark all cycles in $C^{\prime}$. Also, the ILP has a valid solution by, for each cycle $c\in C^{\prime}$, setting $x_{c}$ to the number of times this cycle was removed from $p$. Thus, in this case the algorithm will decide that the given instance is a yes-instance.
Computing $\min(a)$ and $\max(a)$ for all $a\in A$ takes $\mathcal{O}\!\left(|T|+|S|\right)$ time. Combining the choices for the path and $C^{\prime}$ gives at most $2^{k}\cdot|A|^{k}\cdot 2^{2^{k}\cdot|A|^{k}}$ branches. Testing a set $C^{\prime}$ by the marking procedure takes $\mathcal{O}\!\left(|C|^{2}k\right)$ time, but this time is dominated by the time to solve the ILP. The ILP has at most $|C|\leq 2^{k}\cdot|A|^{k}$ variables and $2|A|+|C|={\mathcal{O}\!\left(2^{k}\cdot|A|^{k}\right)}$ constraints, in which the absolute values of all coefficients are bounded by $|T|$.
By the result of Lenstra (1983) and its subsequent improvements Kannan (1987); Frank and Tardos (1987) an ILP instance ${\cal I}_{\text{ILP}}$ with $n$ variables can be solved in $\mathcal{O}\!\left(n^{O(n)}\cdot|{\cal I}_{\text{ILP}}|\right)$ time. We note that, for the given ILP, $|I_{\text{ILP}}|={\mathcal{O}\!\left(2^{k}\cdot|A|^{k}\cdot\log(|T|)\right)}$. Thus, solving the ILP takes $\mathcal{O}\!\left((2^{k}\cdot|A|^{k})^{O(2^{k}\cdot|A|^{k})}\cdot 2^{2k}\cdot|A|^{2k}\cdot\log(|T|)\right)$ time. Hence, the total runtime can be upper-bounded by
$\mathcal{O}\Big{(}2^{k}\cdot|A|^{k}\cdot 2^{2^{k}\cdot|A|^{k}}\cdot(2^{k}\cdot|A|^{k})^{O(2^{k}\cdot|A|^{k})}\cdot 2^{2k}\cdot|A|^{2k}\cdot$
$\log(|T|)+|T|+|S|\Big{)}$
$={\mathcal{O}\!\left(2^{2^{k}\cdot|A|^{k}}\cdot(2|A|)^{O(k\cdot 2^{k}\cdot|A|^{k})}\cdot\log(|T|)+|T|\right)}.$
Recalling the bound on $|A|$ argued in the first paragraph of the proof, we obtain that any instance of Action Executability in which the task network is primitive and $\prec^{+}=\emptyset$ can be solved in time $\mathcal{O}(f(k)\cdot\log(|{\cal I}|)+|{\cal I}|^{3})$, for some computable function $f$ of the number $k$ of states. ∎
Using Generalized Partial Order Width.
At this point, we have seen that all of the considered problems on HTNs become polynomial-time solvable on chains and antichains. In this subsection, we unify and extend these positive results by establishing tractability for all HTNs of constant generalized partial order width (whereas chains and antichains have a generalized partial order width of $1$ and 00, respectively). Intuitively speaking, the following three theorems are obtained by “supercharging” the proof techniques used for chains and antichains via a careful dynamic programming routine.
Theorem 11.
Plan Validation is polynomial-time solvable on primitive task networks of constant generalized partial order width.
Proof.
Let $w$ be the generalized partial order width of the input network, $U$ the set of isolated elements in $T$, and recall that the set $V$ of all remaining elements in $T$ induces a partial order of width at most $w$. First, check whether $p$ is an executable sequence with respect to the preconditions of each action. If not, then return that it is a no-instance. Otherwise, decompose $D_{\prec}[V]$ into $w$ chains $c_{1},\ldots,c_{w}$. This can be done in $\mathcal{O}\!\left(|V|^{2.5}\right)$ time Fellows and McCartin (2003). For $0\leq i\leq|T|$, let $R_{i}[h_{1},\ldots,h_{w}]$ be a boolean variable that is TRUE if and only if the sequence of the first $i$ actions in $p$ is executable in $tn$ such that exactly the first $h_{j}$ tasks are used in each chain $c_{j},j\in[w]$. We compute this variable for all possible assignments in a dynamic programming manner. Initialize all values to FALSE except for $R_{0}[0,\ldots,0]$, which is initialized to TRUE. This completes the calculations for $i=0$. Assume that the calculations for some arbitrary, but fixed $i$ are complete.
Then, update the values for $i+1$ as follows. Iterate over all assignments $R_{i}[h_{1},\ldots,h_{w}]$ that are TRUE. Let $T^{\prime}\subseteq U\cup V$ be such that $t\in T^{\prime}$ if and only if $\alpha(t)=p[i+1]$ and, for each chain $c_{j}$, all the predecessors of $t$ in $\prec^{+}$ in $c_{j}$ are among the first $h_{j}$ tasks in $c_{j}$. Intuitively, $T^{\prime}$ contains the tasks that are executable as the $(i+1)^{\text{th}}$ task in the plan given that the first $h_{j}$ tasks are used in each chain $c_{j}$. Further, let $V^{\prime}\subseteq V$ be such that $t\in V^{\prime}$ if and only if $\alpha(t)=p[i+1]$ and $t$ is among the first $h_{j}$ tasks in its chain $c_{j}$, that is, $V^{\prime}$ contains the already executed tasks of the current action from the chains. Let $o$ denote the number of occurrences of the action $p[i+1]$ in the first $i$ elements of the plan $p$. If $|U\cap T^{\prime}|>o-|V^{\prime}|$, then set $R_{i+1}[h_{1},\ldots,h_{w}]$ to be TRUE. Intuitively, this means that the next task can be taken from $U$ since there is an available task in $U$ that is not already required earlier in the plan. Also, the $(i+1)^{\text{th}}$ task may come from a chain. Thus, for each chain $c_{j},j\in[w],$ such that the $(h_{j}+1)^{\text{th}}$ task is in $T^{\prime}$, set $R_{i+1}[h_{1},\ldots,h_{j-1},h_{j}+1,h_{j+1},\ldots,h_{w}]$ to TRUE. Continue until the computations for $i=|T|$ are completed. Then, the given instance is a yes-instance if and only if there is $R_{|T|}[h_{1},\ldots,h_{w}]$ that is TRUE.
We show by induction over $i$ that the algorithm sets all variables correctly. For $i=0$, the algorithm sets all $R_{0}[h_{1},\ldots,h_{w}]$ correctly as the only way to build an empty plan is to take no tasks from any chain. Now, assume that the algorithm correctly assigned all variables up to some arbitrary, but fixed $i$.
If the algorithm sets some value $R_{i+1}[h_{1},\ldots,h_{w}]$ to TRUE, then we distinguish between two cases. First, assume it does so starting from $R_{i}[h_{1},\ldots,h_{w}]$ being TRUE. Due to the induction hypothesis, there is an executable task sequence that uses the first $h_{j}$ tasks of each chain $c_{j}$, $j\in[w]$, and whose actions correspond to the first $i$ actions of $p$. Action $p[i+1]$ occurs $o$ times among these. As $|V^{\prime}|$ contains all used tasks with this action from the chains, exactly $o-|V^{\prime}|$ executed tasks of this action are in $U$. As $T^{\prime}\cap U$ contains all executable tasks in $U$, if $|T^{\prime}\cap U|>o-|V|$, then there is a task $t\in U$ that is executable for action $p[i+1]$ and has not already been used. Thus, it is correct to set $R_{i+1}[h_{1},\ldots,h_{w}]$ to TRUE. In the other case, there is $j\in[w]$ such that $R_{i}[h_{1},\ldots,h_{j-1},h_{j}-1,h_{j+1},\ldots,h_{w}]$ is TRUE and the $h_{j}^{\text{th}}$ action of the chain $c_{j}$ is executable. Thus, in this case it is also correct to set $R_{i+1}[h_{1},\ldots,h_{w}]$ to TRUE.
For the other direction, suppose there is an executable sequence $r$ of $i$ different tasks in $T$ of the required actions that uses the first $h_{j}$ tasks of each chain $c_{j}$, that is, $R_{i}[h_{1},\ldots,h_{w}]$ should be TRUE. Let the last task in this sequence be $t$ and let $r^{\prime}$ denote the sequence $r$ without $t$. If $t\in U$, then the existence of $r^{\prime}$ implies that $R_{i-1}[h_{1},\ldots,h_{w}]$ is TRUE, and in the computations starting from there we have $|T^{\prime}\cap U|>o-|V|$. If $t\in V$, then $t$ is the $h_{j}^{\text{th}}$ task of a chain $c_{j}$, $R_{i-1}[h_{1},\ldots,h_{j-1},h_{j}-1,h_{j+1},\ldots,h_{w}]$ is TRUE, and in the computations starting from there we have $t\in T^{\prime}$. In both cases, the algorithm correctly sets $R_{i}[h_{1},\ldots,h_{w}]$ to TRUE.
This procedure sets $\mathcal{O}\!\left(|T|\cdot|T|^{w}\right)$ values to TRUE. For each of these, computing $T^{\prime},V^{\prime},$ and $o$ takes $\mathcal{O}\!\left(|T|+|\prec|\right)$ time and it can update at most $w+1$ variables. Hence, setting all variables takes ${\mathcal{O}\!\left(|T|^{w+1}(|T|+|\prec|+w+1)\right)}={\mathcal{O}\!\left(|T|^{w+1}(|T|+|\prec|)\right)}$ time, which dominates the $\mathcal{O}\!\left(|V|^{2.5}\right)$ time it takes to decompose $D_{\prec}[V]$ into $w$ chains. In particular, the running time can be upper-bounded by $|{\cal I}|^{\mathcal{O}(w)}$. ∎
We note that the above algorithm is able to also solve the problem on more general networks, specifically even if the isolated vertices in $U$ are allowed to have in-edges from $V$.
Next, we set our sights on State Reachability. While this problem could be solved via a stand-alone argument on antichains in Theorem 8, in the more general setting considered here, we will first need to apply a reduction rule that can be seen as a simpler version of Reduction Rule 1 used for Action Executability.
Reduction Rule R2:
Assume we are given a primitive State Reachability instance. Let $a_{1}$ and $a_{2}$ be two equivalent actions. For all $t\in T$ such that $\alpha(t)=a_{2}$, set $\alpha(t)=a_{1}$.
Observation 12.
Reduction Rule R2 is safe.
Proof.
Assume there is a task sequence $t_{1},\ldots,t_{\ell}$ solving the original instance. Then, the same sequence solves the reduced instance as it gives the same path in the state transition graph in either instance. The same holds for the other direction. ∎
With this, we are now ready to also establish tractability for State Reachability.
Theorem 13.
State Reachability is polynomial-time solvable on primitive task networks with constantly many states and constant generalized partial order width.
Proof.
Let $k$ be the number of states and $w$ the generalized partial order width of the input instance. Set $U$ to be the set of isolated elements in $T$, and recall that the set $V$ of all remaining elements in $T$ induces a partial order of width at most $w$. We begin by exhaustively applying Reduction Rule R2, which can be done in time at most ${\mathcal{O}\!\left(|{\cal I}|^{3}k^{2}\right)}$. Afterwards, we note that there is at most one action for each of the at most $(k+1)^{k}$ equivalence classes.
The algorithm now decomposes $D_{\prec}[V]$ into $w$ chains $c_{1},\ldots,c_{w}$. This can be done in $\mathcal{O}\!\left(|V|^{2.5}\right)$ time Fellows and McCartin (2003). Let $A={\{a_{1},\ldots,a_{|A|}\}}$. For each state $s$ in the state transition graph, let $R_{s}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{|A|}]$ be a boolean variable that is TRUE if and only if there is an executable task sequence in $tn$ starting from $s_{0}$ and ending in state $s$ that employs exactly the first $h_{j}$ tasks of each chain $c_{j}$, $j\in[w]$, and contains exactly $u_{i}$ tasks with action $a_{i}$, for all $i\in[|A|]$.
We compute this variable for all possible assignments in a dynamic programming manner. Initialize all values to FALSE except for $R_{s_{0}}[0,\ldots,0][0,\ldots,0]$, which is initialized to TRUE. Whenever a variable $R_{s}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{|A|}]$ that was FALSE before is set to TRUE, also perform the following update step. Let $T^{\prime}\subseteq T$ be such that $t\in T^{\prime}$ if, for each chain $c_{j}$, all the predecessors of $t$ in $\prec^{+}$ in $c_{j}$ are among the first $h_{j}$ tasks in $c_{j}$. Consider all the states $s^{\prime}$ that are adjacent to $s$ in ${\cal G}$. If there is $j\in[w]$ such that the $(h_{j}+1)^{\text{th}}$ task of the chain $c_{j}$ is in $T^{\prime}$, has action $a_{i}$, and $a_{i}$ connects $s$ and $s^{\prime}$ in ${\cal G}$, then set $R_{s^{\prime}}[h_{1},\ldots,h_{j-1},h_{j}+1,h_{j+1},\ldots,h_{w}][u_{1},\ldots,u_{i-1},u_{i}+1,u_{i+1},\ldots,u_{|A|}]$ to TRUE. This corresponds to taking the next task from a chain and adding it to a given sequence. Another update rule corresponds to adding a task from $U$ instead. If there is an action $a_{i}$ that connects $s$ and $s^{\prime}$ in ${\cal G}$, then define the following. Let $T_{i}={\{t\in T\mid\alpha(t)=a_{i}\}}$ and let $v_{i}$ denote the number of tasks in $V\cap T_{i}$ that are among the first $h_{j}$ tasks of a chain $c_{j}$. If $u_{i}-v_{i}<|T_{i}\cap T^{\prime}\cap U|$, then set $R_{s^{\prime}}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{i-1},u_{i}+1,u_{i+1},\ldots,u_{|A|}]$ to TRUE. Once no more variables are updated and set to TRUE, the computation is complete. Then, a given State Reachability instance is a yes-instance if and only if there is $R_{s_{g}}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{w}]$ that is TRUE. A given Action Executability instance is a yes-instance if and only if there is $R_{s}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{w}]$ that is TRUE such that, for each action $a_{i}\in S$, we have $\min(a)\geq u_{i}$.
We show that the algorithm sets all variables correctly by an inductive argument over $\sum_{i=1}^{|A|}u_{i}$. Note that when one TRUE variable sets another variable to TRUE, this increases this sum by exactly 1. For $\sum_{i=1}^{|A|}u_{i}=0$, the algorithm sets all variables correctly as the only way to build an empty plan is to take no tasks from any chain and remain in the initial state, and thus, the only TRUE variable in this case is $R_{s_{0}}[0,\ldots,0][0,\ldots,0]$. Now, assume that the algorithm correctly assigned all variables up to some arbitrary, but fixed $\sum_{i=1}^{|A|}u_{i}$.
If some value $R_{s}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{w}]$ is set to TRUE by the algorithm, then we distinguish between two cases. First, assume it does so starting from some $R_{s^{\prime}}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{i-1},u_{i}-1,u_{i+1},\ldots,u_{|A|}]$ being TRUE with $i\in[|A|]$. Then, $a_{i}$ is executable in state $s^{\prime}$ and transforms it into $s$. Further, due to the induction hypothesis, there is an executable task sequence that uses the first $h_{j}$ tasks of each chain $c_{j}$, $j\in[w]$, and consists of $u_{\ell}$ tasks of action $a_{\ell}$, for each $\ell\in[|A|]\setminus{\{i\}}$, and $u_{i}-1$ tasks of action $a_{i}$. Thus, it contains $u_{i}-v_{i}$ tasks in $U$ with action $a_{i}$. There are more than $u_{i}-v_{i}$ executable tasks of action $a_{i}$ in $U$, and thus, one of the remaining tasks of action $a_{i}$ can be appended to the sequence. The resulting sequence is executable in $tn$ starting from $s_{0}$, uses the same number of tasks of each chain and action, except one more task with action $a_{i}$. Thus, it is correct to set $R_{s^{\prime}}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{w}]$ to TRUE. In the other case, it sets the variable to TRUE starting from some $R_{s^{\prime}}[h_{1},\ldots,h_{j-1},h_{j}-1,h_{j+1},h_{w}][u_{1},\ldots,u_{i-1},u_{i}-1,u_{i+1},\ldots,u_{|A|}]$ being TRUE with $j\in[w]$, $i\in[|A|]$. Then, due to the induction hypothesis, there is an executable task sequence that uses the first $h_{\ell^{\prime}}$ tasks of each chain $c_{\ell^{\prime}}$, for each $\ell^{\prime}\in[w]\setminus\{j\}$, but only uses the first $h_{j}-1$ tasks of the chain $c_{j}$, and it consists of $u_{\ell}$ tasks of action $a_{\ell}$, for each $\ell\in[|A|]\setminus{\{i\}}$, and $u_{i}-1$ tasks of action $a_{i}$. Further, the $h_{j}^{\text{th}}$ task of chain $h_{j}$ has action $a_{i}$, is executable at the end of this sequence, and transforms state $s^{\prime}$ into $s$. Thus, in this case it is also correct to set the variable to TRUE.
For the other direction, suppose there is an executable sequence $r$ of $\sum_{i=1}^{|A|}u_{i}$ different tasks in $T$ such that, for each $\ell\in[|A|]$, there are $u_{\ell}$ tasks with action $a_{\ell}$ that use the first $h_{j}$ tasks of each chain $c_{j}$. Suppose $r$ ends in state $s$, that is, $R_{s}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{|A|}]$ should be TRUE. Let the last task in this sequence be $t$ with $\alpha(t)=a_{i}$ and let $r^{\prime}$ denote the sequence $r$ without $t$. Let $r^{\prime}$ end in state $s^{\prime}$. If $t\in U$, then the existence of $r^{\prime}$ implies $R_{s^{\prime}}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{i-1},u_{i}-1,u_{i+1},\ldots,u_{|A|}]$ is TRUE. In the update step of this variable, $t$ is executable, and thus, $u_{i}-v_{i}<|T_{i}\cap T^{\prime}\cap U|$. Then, the algorithm correctly sets $R_{s}[h_{1},\ldots,h_{w}][u_{1},\ldots,u_{|A|}]$ to TRUE. If $t\in V$, then $t$ is the $h_{j}^{\text{th}}$ task of a chain $c_{j}$, $R_{s^{\prime}}[h_{1},\ldots,h_{j-1},h_{j}-1,h_{j+1},\ldots,h_{w}][u_{1},\ldots,u_{i-1},u_{i}-1,u_{i+1},\ldots,u_{|A|}]$ is TRUE, and in the computations starting from there $t\in T^{\prime}$. In both cases, the algorithm correctly sets the variable to TRUE.
For each variable that is TRUE, we have that $s$ is among the $k$ states in ${\cal G}$, $0\leq h_{j}\leq|T|$ for $j\in[w]$, and $0\leq\sum_{i=1}^{|A|}u_{i}\leq|T|$. Thus, at most ${\mathcal{O}\!\left(|T|^{w}\cdot k\cdot|T|^{|A|}\right)}$ variables are considered. For each of these, computing $T^{\prime}$, the $u_{i}$, and $v_{i}$ takes $\mathcal{O}\!\left(|T|+|\prec|\right)$ time, it can update at most $(w+1)\cdot|A|$ variables, and it can be checked in $\mathcal{O}\!\left(|A|\right)$ time if it yields a yes-instance. Hence, the algorithm takes ${\mathcal{O}\!\left(|T|^{w+|A|}\cdot k\cdot(|T|+|\prec|+(w+1)\cdot|A|)\right)}={\mathcal{O}\!\left(|T|^{w+|A|}\cdot k\cdot(|T|+|\prec|+w|A|)\right)}$ time. Once again, the theorem now follows by the bound on $|A|$ established in the first paragraph of the proof. ∎
We observe that, once more, the above algorithm is actually able to solve the more general problem where the isolated vertices in $U$ are allowed to have in-edges from $V$. Finally, we note that Reduction Rule R1 by itself does not suffice to transfer this result to Action Executability as the number of neighborhoods of tasks is not bounded (as it was for Theorem 10). Nevertheless, the desired result can be obtained by carefully adapting the dynamic program from Theorem 13 and incorporating the idea of the reduction rule on the (isolated) tasks in $U$.
Theorem 14.
Action Executability is polynomial-time solvable on primitive task networks with constantly many states and constant generalized partial order width.
Proof.
Let $k$ be the number of states and $w$ the generalized partial order width of the input instance. Set $U$ to be the set of isolated elements in $T$, and recall that the set $V$ of all remaining elements in $T$ induces a partial order of width at most $w$. The algorithm is similar to the one presented for Theorem 13. First, test whether Reduction Rule R0 can be applied, and if yes we use it to terminate. Otherwise, decompose $D_{\prec}[V]$ into $w$ chains $c_{1},\ldots,c_{w}$. This can be done in $\mathcal{O}\!\left(|V|^{2.5}\right)$ time Fellows and McCartin (2003). Let $A={\{a_{1},\ldots,a_{|A|}\}}$. Let $E={\{e_{1},\ldots,e_{|E|}\}}$ denote the set of equivalence classes among all actions. For each state $s$ in the state transition graph, let $R_{s}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{|E|}]$ be a boolean variable that is TRUE if and only if there is an executable task sequence in $tn$ starting from $s_{0}$ and ending in state $s$ that employs exactly the first $h_{j}$ tasks of each chain $c_{j}$, $j\in[w]$, and contains exactly $r_{i}$ tasks from $U$ with action equivalence class $e_{i}$, for all $i\in[|E|]$.
We compute this variable for all possible assignments in a dynamic programming manner. Initialize all values to FALSE except for $R_{s_{0}}[0,\ldots,0][0,\ldots,0]$, which is initialized to TRUE. Whenever a variable $R_{s}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{|E|}]$ that was FALSE before is set to TRUE, also perform the following update step. Let $T^{\prime}\subseteq T$ be such that $t\in T^{\prime}$ if, for each chain $c_{j}$, all the predecessors of $t$ in $\prec^{+}$ in $c_{j}$ are among the first $h_{j}$ tasks in $c_{j}$. Consider all the states $s^{\prime}$ that are adjacent to $s$ in ${\cal G}$. If there is $j\in[w]$ such that the $(h_{j}+1)^{\text{th}}$ task of the chain $c_{j}$ is in $T^{\prime}$, has action $a_{i}$, and $a_{i}$ connects $s$ and $s^{\prime}$ in ${\cal G}$, then set $R_{s^{\prime}}[h_{1},\ldots,h_{j-1},h_{j}+1,h_{j+1},\ldots,h_{w}][r_{1},\ldots,r_{|E|}]$ to TRUE. This corresponds to taking the next task from a chain and adding it to a given sequence. Another update rule corresponds to adding a task from $U$ instead. If there is an action equivalence class $e_{i}$ that connects $s$ and $s^{\prime}$ in ${\cal G}$, then define the following. Let $T_{i}={\{t\in T\mid\alpha(t)\in e_{i}\}}$ and let $v_{i}$ denote the number of tasks in $V\cap T_{i}$ that are among the first $h_{j}$ tasks of a chain $c_{j}$. If $r_{i}-v_{i}<|T_{i}\cap U|$, then set $R_{s^{\prime}}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{i-1},r_{i}+1,r_{i+1},\ldots,r_{|E|}]$ to TRUE. Once no more variables are updated and set to TRUE, the computation is complete.
The algorithm sets all variables correctly by the same reasoning as in the proof of Theorem 13. For any $R_{s}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{|E|}]$ and action $a\in A$, let $c(a)$ denote the number of tasks in $V$ with action $a$ that are among the first $h_{j}$ tasks of a chain $c_{j}$. Then, the algorithm returns that it is a yes-instance if and only if there is $R_{s}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{|E|}]$ that is TRUE such that, for each action equivalence class $e_{i}\in E$, we have $\sum_{a\in e_{i}}\max{\{0,\min(a)-c(a)\}}\leq r_{i}$, and, for each action $a\in A$, there are at least $\min(a)-c(a)$ tasks of action $a$ in $U$.
Suppose there is such a TRUE variable. Then, there is an executable task sequence that uses $c(a)$ tasks in $V$ with action $a$ for each $a\in A$, and uses $r_{i}$ tasks from $U$ with action in equivalence class $e_{i}$ for all $i\in[|E|]$. As all tasks in $U$ are isolated, two tasks from $U$ can be used interchangeably in a task sequence if they are from the same action equivalence class. Tasks in the sequence can be replaced by tasks from $U$ of the same equivalence class such that there are at least $\min(a)-c(a)$ tasks with action $a$ from $U$ in the sequence for all $a\in A$. Then, the sequence is a solution to Action Executability.
For the other direction, suppose there is a task sequence that solves the Action Executability instance. Let it result in some state $s$, let it use the first $h_{j}$ tasks of each chain $c_{j}$, and let it employ $r_{i}$ tasks of equivalence class $e_{i}$ for all $i\in[|E|]$. Then, the variable $R_{s}[h_{1},\ldots,h_{w}][r_{1},\ldots,r_{|E|}]$ is TRUE. Further, as the sequence contains at least $\min(a)$ tasks of each action $a\in A$, and only $c(a)$ of these are tasks in $V$, it contains at least $\min(a)-c(a)$ tasks of action $a$ from $U$. In particular, this also gives $\sum_{a\in e_{i}}(\max{\{0,\min(a)-c(a)\}})\leq r_{i}$, so the algorithm correctly returns that it is a yes-instance.
Computing the equivalence classes takes $\mathcal{O}\!\left(|{\cal I}|^{3}\right)$ time. As there are $|E|\leq(k+1)^{k}$ equivalence classes, at most $\mathcal{O}\!\left(|T|^{w}\cdot k\cdot|T|^{(k+1)^{k}}\right)$ variables are considered. For each of these, computing all $v_{i}$ takes $\mathcal{O}\!\left(|T|\right)$ time by iterating through all chains time and it can update at most $w+|E|$ variables. Further, it takes $\mathcal{O}\!\left(|T|\right)$ time to check whether a variable yields a yes-instance. Hence, the algorithm takes ${\mathcal{O}\!\left(|T|^{w+(k+1)^{k}}\cdot k\cdot(|T||+w+|E|)+|V|^{2.5}+|\prec|\right)}={\mathcal{O}\!\left(|T|^{w+(k+1)^{k}+1}\cdot k\right)}$ time. ∎
We recall that Theorem 14 also immediately implies the analogous result for Plan Existence.
5 Hierarchical Task Networks
Considering HTNs with compound tasks adds another, seemingly opaque layer of difficulty to the considered planning problems. In this section, we show that—at least in terms of classical complexity—we can cleanly characterize the jump from primitive to compound through three measures on compound task networks: we obtain an algorithmic meta-theorem that allows to lift polynomial-time solvability from primitive to compound task networks, if all three of these measures are bounded. Moreover, this result is tight in the sense that none of the three measures can be dropped from the meta-theorem.
The three measures we need are ${C_{\#}}$, ${C_{s}}$, and ${C_{d}}$:
Definition 5.1.
For an HTN planning problem with initial task network $tn=(T,\prec^{+},\alpha)$, we refer to the number of compound tasks in $tn$ by ${C_{\#}}=|{\{t\in T\mid\alpha(t)\in{\cal C}\}}|$. The maximum size of a network that a task can be decomposed into is ${C_{s}}=\max{\{|T_{m}|\mid(c,(T_{m},\prec^{+}_{m},\alpha_{m}))\in M\}}$. The maximum depth $C_{d}$ of a network is defined recursively: ${C_{d}}=0$ for primitive task networks, and a compound task network has ${C_{d}}=i$ if $i$ is the smallest integer such that every decomposition of all compound tasks in the initial network $T$ results in a new network where ${C_{d}}\leq i-1$.
For our later considerations, it will be useful to additionally define the maximum number of pairwise non-isomorphic networks that a task can be decomposed into; we denote that by ${C_{c}}=\max_{c\in{\cal C}}|{\{tn_{m}\mid(c,tn_{m})\in M\}}|$. The next technical lemma shows how these measures help us bound the “decomposition complexity” of compound HTNs.
Lemma 15.
A task network $tn=(T,\prec^{+},\alpha)$ can be decomposed into at most ${C_{c}}^{\sum_{i=0}^{{C_{d}}-1}{C_{\#}}\cdot{C_{s}}^{i}}$ pairwise non-isomorphic primitive task networks $tn^{\prime}=(T^{\prime},{{\prec^{+}}}^{\prime},\alpha^{\prime})$. For each of these, $|T^{\prime}|\leq|T|+{C_{\#}}\cdot({C_{s}}^{{C_{d}}}-1)$.
Proof.
We first show the second part of the statement. The $|T|-{C_{\#}}$ primitive tasks in $tn$ are still present in any network $tn^{\prime}$ that $tn$ can be decomposed into. A compound task in $tn$ can be decomposed into at most $C_{s}$ compound tasks in one decomposition step and there are at most ${C_{d}}$ decomposition steps. Thus, each compound task in $tn$ can create at most ${C_{s}}^{C_{d}}$ tasks in $tn^{\prime}$. Hence, $|T^{\prime}|\leq|T|-{C_{\#}}+{C_{\#}}\cdot{C_{s}}^{C_{d}}$.
For the first statement, note that decomposing each of the ${C_{\#}}$ compound tasks once in a task network of depth ${C_{d}}$ can result in at most ${C_{c}}^{C_{\#}}$ pairwise non-isomorphic task networks of depth at most ${C_{d}}-1$. Further, each of these resulting networks has at most ${C_{\#}}\cdot{C_{s}}$ compound tasks. Thus, $tn$ can be decomposed into at most
${C_{c}}^{{C_{\#}}}\cdot{C_{c}}^{{C_{\#}}\cdot{C_{s}}}\cdot\ldots\cdot{C_{c}}^{{C_{\#}}\cdot{C_{s}}^{{C_{d}}-1}}={C_{c}}^{\sum_{i=0}^{{C_{d}}-1}{C_{\#}}\cdot{C_{s}}^{i}}$
pairwise non-isomorphic task networks. ∎
Having defined our measures of interest, we proceed to formalize the set of problems our meta-theorem will capture. Intuitively, the aim here is to formalize the set of all problems which deal with compound HTNs by decomposing them.
Definition 5.2.
Let PR be an arbitrary HTN planning problem. Then, PR is decomposable if and only if, for all non-primitive task networks $tn$, PR on $tn$ is a yes-instance if and only if there is a primitive task network $tn^{\prime}$ such that $tn$ can be decomposed into $tn^{\prime}$ and $tn^{\prime}$ is a yes-instance of PR.
Note that all problems discussed in this work are decomposable. Next, in order to capture the general notions of “tractability” for various HTN problems on networks where certain measures are bounded, we need to slightly restrict the measures in question, to avoid entirely degenerate measures.
Definition 5.3.
Let $\kappa$ be any numerical measure of HTNs, i.e., a mapping that assigns each HTN a non-negative number. We say that $\kappa$ is stable if there exists a computable function $f$ with the following property: for each primitive HTN $tn$ and each primitive HTN $tn^{\prime}$ obtained from $tn$ by adding or removing a single task, $\kappa(tn^{\prime})\leq f(\kappa(tn))$.
In particular, all measures considered here are stable. The number of states in the state transition graph only depends on the initial state $s_{0}$ and the set of actions $A$, so it does not change when applying decomposition methods. Moreover, adding an element into the network can only increase the generalized partial order width (as well as all the other structural measures mentioned in the manuscript) by at most $1$.
Theorem 16.
Let PR be a decomposable HTN planning problem and $\kappa$ a stable measure. Assume PR is polynomial-time solvable when restricted to primitive networks of constant $\kappa$. Then, PR is polynomial-time solvable when restricted to networks of constant $\kappa$, ${C_{d}},{C_{\#}},{C_{s}}$.
Proof.
Let ${\cal I}$ have task network $tn=(T,\prec^{+},\alpha)$. If there are no compound tasks in $T$, i.e., ${C_{d}}=0$, the statement immediately holds by applying an algorithm for primitive networks. Thus, assume ${C_{d}}\geq 1$. Then, PR in $tn$ can be solved by listing all the primitive instances $tn$ can be decomposed into and solving PR in each of these. By Lemma 15, there are at most ${C_{c}}^{g({C_{d}},{C_{\#}},{C_{s}})}\leq|{\cal I}|^{g({C_{d}},{C_{\#}},{C_{s}})}$ such instances ${\cal I}_{p}$ for $g({C_{d}},{C_{\#}},{C_{s}})=\sum_{i=0}^{{C_{d}}-1}{C_{\#}}\cdot{C_{s}}^{i}$ and, for each of these, $|{\cal I}_{p}|\leq|{\cal I}|+f^{\prime\prime}({C_{d}},{C_{\#}},{C_{s}})$ for some computable function $f^{\prime\prime}$. Further, for each task network $tn_{p}$ of an instance ${\cal I}_{p}$, we have $\kappa(tn_{p})\leq f^{\prime\prime\prime}(\kappa,{C_{d}},{C_{\#}},{C_{s}})$, for some computable function $f^{\prime\prime\prime}$ as $\kappa$ is stable and $tn_{p}$ is obtained from $tn$ by removing ${C_{\#}}$ compound tasks and adding $f^{\prime\prime}({C_{d}},{C_{\#}},{C_{s}})$ primitive tasks. Thus, if primitive instances ${\cal I}_{p}$ can be solved in time at most $|{\cal I}_{p}|^{f(\kappa(tn_{p}))}$, then the total runtime is
$\mathcal{O}\Big{(}(|{\cal I}|+f^{\prime\prime}({C_{d}},{C_{\#}},{C_{s}}))^{f(f^{\prime\prime\prime}(\kappa,{C_{d}},{C_{\#}},{C_{s}}))}$
$\cdot{C_{c}}^{g({C_{d}},{C_{\#}},{C_{s}})}\Big{)}$
$={\mathcal{O}\!\left(|{\cal I}|^{f^{\prime}(\kappa,{C_{d}},{C_{\#}},{C_{s}})}\right)},$
where $f^{\prime}(\kappa,{C_{d}},{C_{\#}},{C_{s}})=g({C_{d}},{C_{\#}},{C_{s}})+\\ f(f^{\prime\prime\prime}(\kappa,{C_{d}},{C_{\#}},{C_{s}}))+f^{\prime\prime}({C_{d}},{C_{\#}},{C_{s}})^{f(f^{\prime\prime\prime}(\kappa,{C_{d}},{C_{\#}},{C_{s}}))}$. ∎
We complement Theorem 16 by providing lower bounds which show that bounding $C_{d}$, $C_{\#}$, and $C_{s}$ is necessary in order to obtain the result. In particular, we show that if any one of these measures is not bounded by a constant, then some of the problems considered in this article become NP-hard on compound networks, even for settings that were easily solvable on primitive networks. We note that there are several other existing reductions showing the hardness that derives from decomposing non-primitive networks, see, e.g., Erol et al. (1996); Behnke et al. (2015). However, our reductions are stronger in the sense that they bound all but one of the parameters $C_{d}$, $C_{\#}$, $C_{s}$.
Recall that, by Theorem 11, Plan Validation is solvable in polynomial time on primitive networks of constant generalized partial order width. Moreover, as per Theorem 10, the same also holds for Action Executability when additionally bounding the number of states. In contrast, in Theorems 17, 18 and 19, we prove Plan Validation (and also Action Executability) to be NP-hard on general networks even if:
  • any two of $C_{d}$, $C_{\#}$, $C_{s}$, are constant,
  • all possible decompositions of the network have a generalized partial order width of 00,
  • there is only a single state, and
  • $C_{c}$ is also constant.
Theorem 17.
Plan Validation and Action Executability are NP-hard even if ${C_{c}}=7$, ${C_{d}}=1$, ${C_{s}}=2$, $\prec^{+}=\emptyset$, and $\operatorname{prec}(a)$=$\operatorname{del}(a)$=$\operatorname{add}(a)$$\emptyset$ for every $a\in A$.
Proof.
We prove hardness by reducing from 3-SAT-(2,2), an NP-hard 3-SAT variant, where among all clauses each variable appears exactly 4 times: 2 times as a positive literal and 2 times as a negated literal Darmann and Döcker (2021).
We construct an HTN planning instance where the set of actions $A={\{x_{1},\overline{x_{1}},\ldots,x_{n},\overline{x_{n}}\}}$ contains an action for each literal. The initial task network contains a compound task $t_{x}$ for each variable $x$ and a compound task $t_{c}$ for each clause. A variable compound task $t_{x}$ can be decomposed into either a task network containing two tasks with action $x$ or a task network containing two tasks with action $\overline{x}$. Decomposing $t_{x}$ into the first (second, respectively) one of these is interpreted as setting $x$ to TRUE (FALSE, respectively). A clause compound task $t_{c}$ where the clause $c$ contains literals $\ell_{1},\ell_{2},\ell_{3}$ can be decomposed into one of 7 subnetworks, one with no tasks at all, one with a single task with action $\ell_{1},\ell_{2},$ or $\ell_{3}$, or one with two tasks with actions $\ell_{1}$ and $\ell_{2}$, actions $\ell_{1}$ and $\ell_{3}$, or actions $\ell_{2}$ and $\ell_{3}$. Decomposing a clause task is interpreted as acknowledging that exactly the described literals are FALSE. This gives an instance with $\prec^{+}=\emptyset$. For Plan Validation, let the input plan $p$ be any sequence that contains each literal exactly twice. For Action Executability, let $S$ contain each literal exactly twice.
If we are given a yes-instance of 3-SAT-(2,2), consider a satisfying variable assignment and decompose the compound tasks accordingly using the interpretations given above. Then, in the resulting primitive task network, each literal appears exactly twice as follows. If a variable $x$ is set to TRUE, then $t_{x}$ creates exactly 2 tasks with action $x$ and there is no clause task creating a task with action $x$. At the same time, the literal $\overline{x}$ appears in exactly 2 clauses and, as it is FALSE, each of these clause tasks creates a task with action $\overline{x}$. The case in which the variable $x$ is set to FALSE is analogous. As the tasks are executable in any order, this yields a yes-instance for Plan Validation and Action Executability.
For the other direction, suppose we have a yes-instance for Plan Validation or Action Executability. Then, the constructed HTN can be decomposed into a primitive task network with an executable plan such that each literal appears at least twice. Consider the variable assignment given by interpreting the decomposition of the variable tasks. Towards a contradiction, assume that there is a clause $c$ with literals $\ell_{1},\ell_{2},\ell_{3}$ that is not satisfied by the assignment. Then, none of these literals is represented by a task decomposed from a variable task. Further, the clause task $c$ can only be decomposed into at most 2 tasks with different actions from $\ell_{1}$, $\ell_{2}$, and $\ell_{3}$. Without loss of generality, let the decomposition of this clause task not produce a task with action $\ell_{1}$. This literal occurs in exactly one other clause and the corresponding clause task can produce at most one task with action $\ell_{1}$. This contradicts having 2 occurrences of action $\ell_{1}$ in the plan. Thus, all the clauses are satisfied.
Observe that the created task network has depth ${C_{d}}=1$ and each compound task can be decomposed into one of at most ${C_{c}}=7$ different task networks which each contain at most ${C_{s}}=2$ tasks. ∎
Theorem 18.
Plan Validation and Action Executability are NP-hard even if ${C_{c}}=7$, ${C_{d}}=2$, ${C_{\#}}=1$, $\prec^{+}=\emptyset$, and $\operatorname{prec}(a)$=$\operatorname{del}(a)$=$\operatorname{add}(a)$$\emptyset$ for every $a\in A$.
Proof.
Consider the task network $tn$ constructed in the proof of Theorem 17. Build a new initial task network $tn^{\prime}=(T={\{t\}},\prec^{+}=\emptyset,\alpha)$, and let $(\alpha(t),tn)\in M$. Then, ${C_{\#}}=1$, ${C_{d}}=2$, ${C_{c}}=7$, and after decomposing $t$ into $tn$, the reduction works exactly as in the proof of Theorem 17. ∎
Theorem 19.
Plan Validation and Action Executability are NP-hard even if ${C_{c}}=7$, ${C_{\#}}=1$, ${C_{s}}=3$, $\prec^{+}=\emptyset$, and $\operatorname{prec}(a)$=$\operatorname{del}(a)$=$\operatorname{add}(a)$$\emptyset$ for every $a\in A$.
Proof.
We adapt our construction used in the proof of Theorem 17. Let the 3-SAT-(2,2) instance have variables $x_{1},\ldots,x_{n}$ and clauses $C_{1},\ldots,C_{m}$. Once more, let the set of actions $A={\{x_{1},\overline{x_{1}},\ldots,x_{n},\overline{x_{n}}\}}$ contain an action for each literal, and let there be a compound task for each variable and clause, i.e., ${\cal C}={\{X_{1},\ldots,X_{n},C_{1},\ldots,C_{m}\}}$. For each $i\in[n-1]$, the variable task $X_{i}$ can be decomposed into one of two subnetworks. Both these subnetworks contain the compound task $X_{i+1}$, and one of them contains two tasks with action $x_{i}$, while the other contains two tasks with action $\overline{x_{i}}$. The decomposition for $X_{n}$ is similar, but instead of a compound task $X_{n+1}$, it adds the compound task $C_{1}$ in both subnetworks. Just like before, the compound tasks for clauses can be decomposed into one of the 7 subnetworks representing proper subsets of the literals in the clause. For $i\in[m-1]$, they additionally contain the compound task $C_{i+1}$. The initial network now contains only a single compound task $X_{1}$, giving ${C_{\#}}=1$. As before, each compound task can be decomposed into one of at most ${C_{c}}=7$ different task networks which, in this case, each contain at most ${C_{s}}=3$ tasks. However, the decomposition depth is no longer bounded by a constant. For Plan Validation, let the input plan $p$ be any sequence that contains each literal exactly twice. For Action Executability, let $S$ contain each literal exactly twice. The correctness of the reduction follows from the same arguments that were used in the proof of Theorem 17. ∎
We conclude this section with an analogous intractability result complementing Theorem 16, even if we consider the “classical” partial order width as our stable measure, as opposed to the generalized variant considered here:
Theorem 20.
Plan Existence, Action Executability, and State Reachability are NP-hard in task networks where $\prec^{+}$ describes a total order on $T$ and all subnetworks in $M$ are total orders even if ${C_{c}}=2$, ${C_{d}}=1$, and ${C_{s}}=1$. The same holds if ${C_{c}}=2$, ${C_{d}}=2$, and ${C_{\#}}=1$, or ${C_{c}}=2$, ${C_{\#}}=1$, and ${C_{s}}=2$.
Proof.
We reduce from $k$-Multicolored-Clique on a graph $G=(V,E)$, which is NP-hard for unbounded $k$ Fellows et al. (2009). Let $tn(p,d,a)$ with $p,d,a\subseteq F$ refer to a task network with a single task $t$ with action $\alpha(t)$ such that $\operatorname{prec}(\alpha(t))=p$, $\operatorname{del}(\alpha(t))=d$, and $\operatorname{add}(\alpha(t))=a$. For each vertex $v_{i}$, $i\in|V|$, create a compound task $V_{i}$, and, for each edge $e_{i}$, $i\in|E|$, create a compound task $E_{i}$. Create a primitive task $t_{g}$ with action $a_{g}$ such that $\operatorname{prec}(a_{g})={\{\text{EDGE}_{j,j^{\prime}}\mid j,j^{\prime}\in[k],j<j^{\prime}\}}$, $\operatorname{del}(a_{g})=\emptyset$, and $\operatorname{add}(a_{g})={\{\text{GOAL}\}}$. The initial network $tn$ contains exactly these tasks in a total order where $(V_{i},V_{i+1})\in\prec$ for all $i\in[|V|-1]$, $(V_{|V|},E_{1})\in\prec$, $(E_{i},E_{i+1})\in\prec$ for all $i\in[|E|-1]$, and $(E_{|E|},t_{g})\in\prec$. Each vertex and edge compound task can be decomposed into a primitive task with no effect, that is, $(V_{i},tn(\emptyset,\emptyset,\emptyset))\in M$ for all $i\in[|V|]$, and $(E_{i},tn(\emptyset,\emptyset,\emptyset))\in M$ for all $i\in[|E|]$. This corresponds to not selecting respective vertices and edges as part of the clique. Further, a vertex can be chosen for color $j$ if no other vertex has been chosen for this color (represented by having $\text{COLOR}_{j}$ in the state). Thus, $(V_{i},tn({\{\text{COLOR}_{j}\}},{\{\text{COLOR}_{j}\}},{\{v_{i}\}}))\in M$ for $i\in[|V|]$ and the color $j$ of the vertex $v_{i}$. An edge can be chosen as the $j^{\text{th}}$ edge in the clique if both its vertices are chosen. Thus, for all $i\in[|E|]$, if $e_{i}=(v,v^{\prime})$, $v$ has color $j$, and $v^{\prime}$ has color $j^{\prime}$, then $(E_{i},tn({\{v,v^{\prime}\}},\emptyset,{\{\text{EDGE}_{j,j^{\prime}}\}}))\in M$. The construction is concluded by setting the initial state $s_{0}={\{\text{COLOR}_{i}\mid i\in[k]\}}$, setting $S={\{a_{g}\}}$ for Action Executability, and setting $s_{g}={\{\text{GOAL}\}}$ for State Reachability. The construction clearly takes polynomial time.
Suppose there is a multicolored clique of size $k$. Decompose according to the above interpretation. We obtain a primitive network with a total order on the tasks. The corresponding task sequence is executable in $tn$ from $s_{0}$, and as it executes all tasks including $t_{g}$, it yields a yes-instance for Plan Existence, Action Executability, and State Reachability.
For the other direction, suppose any one of the three HTN instances has a solution. This implies that the task $t_{g}$ can be executed. The precondition of $t_{g}$ ensures that an edge is selected for each pair of colors. The preconditions on the edge tasks ensure that all the vertices incident to the selected edges have been selected. The preconditions on the vertex tasks ensure that at most 1 vertex of each color is selected. Thus, we have a yes-instance of $k$-Multicolored-Clique.
For the constructed instances, ${C_{d}}={C_{s}}=1$ and ${C_{c}}=2$. For the two other combinations of parameters for which the statement holds, we employ the same approaches we used to adapt Theorem 17. In particular, to prove the statement for ${C_{c}}=2$, ${C_{d}}=2$, and ${C_{\#}}=1$, we let the initial network contain a single compound task that is decomposed into a network with the above structure. To prove the statement for ${C_{c}}=2$, ${C_{\#}}=1$, and ${C_{s}}=2$, we let the initial network only contain the compound task $V_{1}$. Then, we adapt all the decomposition methods. For each method $(c,tn_{m})\in M$, the network $tn_{m}$ consisted of a single primitive task $t_{m}$. Let $t^{\prime}$ denote the task identifier that followed $c$ in the total order in the initial network. Instead of decomposing into a single primitive task $t_{m}$, let the method now decompose into a network with tasks $t_{m}$ and $t^{\prime}$ with $(t_{m},t^{\prime})\in\prec^{+}$. Note that, in this construction, exactly the same set of primitive networks can be created as in the original construction, but now ${C_{\#}}=1$ and ${C_{d}}$ is not bounded. ∎
6 A Parameterized Analysis of HTNs
In this final section, we ask whether—and under which parameterizations—the fundamental problems on HTNs admit fixed-parameter algorithms.
We begin by noting that while there is by now a very broad range of graph-theoretic parameters which can typically be used to achieve fixed-parameter tractability, none of the “usual suspects” in that regard will help when dealing with any of the four problems of interest here. In particular, Theorems 2 and 5 immediately exclude the algorithmic application of not only treewidth Robertson and Seymour (1986)—the by far most commonly used structural graph parameter—but also treedepth Nesetril and de Mendez (2012), the feedback edge number Ganian and Korchemna (2021) and even the broad family of analogous measures defined specifically on directed graphs Ganian et al. (2016).
Moreover, while the positive results obtained in Section 4 establish that our problems of interest are polynomial-time solvable when the generalized partial order width is fixed, none of the problems is fixed-parameter tractable when parameterized by the generalized partial order width (or even by the “usual” partial order width). Indeed, since the Shuffle Product problem used in the reductions underlying the proofs of Theorems 3 and 4 is also known to be W[2]-hard when parameterized by the width $w$ (van Bevern et al., 2016, Theorem 2.4), it turns out that the proofs of these two theorems also immediately establish the following:
Corollary 21.
Plan Validation on primitive networks where $\prec^{+}$ partitions $T$ into $w$ chains such that there are no dependencies between the chains is W[2]-hard parameterized by $w$, even if $|A|=2$ and $\operatorname{prec}(a)=\operatorname{del}(a)=\operatorname{add}(a)=\emptyset$ for all actions $a\in A$.
Corollary 22.
State Reachability and Plan Existence on primitive networks where $\prec^{+}$ partitions $T$ into $w$ chains are W[2]-hard parameterized by $w$, even if $|A|=5$ and the state transition graph has $k=4$ states.
Fixed-Parameter Tractability via Vertex Cover.
In this subsection, we contrast the above lower bounds with a fixed-parameter algorithm that exploits a different parameterization of the network: we show that Plan Validation is fixed-parameter tractable when parameterized by the vertex cover number of $G_{\prec}$. Recall that a vertex cover of $G_{\prec}$ is a subset of tasks $V\subseteq T$ such that, for all $t\prec t^{\prime}$, $t\in V$ and/or $t^{\prime}\in V$; the vertex cover number is the minimum size of a vertex cover of $G_{\prec}$. We remark that, while the vertex cover number has also been used to obtain fixed-parameter algorithms for many other problems Bhore et al. (2020); Balko et al. (2022); Blazej et al. (2023), the techniques from previous works do not easily translate to HTN planning problems. Instead, our algorithms make use of a delicate branching routine whose correctness requires a careful analysis of the state space.
Theorem 23.
Plan Validation is fixed-parameter tractable parameterized by the vertex cover number of $G_{\prec}$.
Proof.
By definition, $p$ is executable if and only if the actions in the plan have their preconditions fulfilled and each $t\in T$ can be assigned a unique index in the sequence such that the assignment according to $\alpha$ matches the plan and the imposed order does not conflict with $\prec^{+}$. Observe that simulating the action sequence suffices to check whether all preconditions are met at the respective step, so it suffices to handle the assignment of tasks to indices. We branch over all possible orderings of the vertices in $V$. Orderings that do not respect an edge in $\prec^{+}\cap(V\times V)$ are discarded immediately. For $i\in[|V|]$, let $v_{i}$ denote the $i^{\text{th}}$ task in the ordering. Then, for all $t\in T\setminus V$, let the priority of $t$ be defined as $\psi(t)=\min_{(t,v_{i})\in\prec}i$ and $\psi(t)=\infty$ if there is no such edge. For all $v_{i}\in V$, let $\psi(v_{i})=i+0.5$. We now iterate through $p$ and assign tasks to each index in a greedy manner. Assume we have already assigned tasks to the first $i-1$ actions for some $i\in[|T|]$. Then, we say that $t\in T$ is executable as the $i^{\text{th}}$ action if $\alpha(t)=p[i]$, it is not yet assigned an index, and all its predecessors in $\prec^{+}$ are assigned to indices less than $i$. Among all executable tasks at the index $i$, we choose the one with the lowest priority according to $\psi$. If there is no executable task for some index, then we return that this branch does not yield a yes-instance. The algorithm returns that it is a yes-instance if any branch manages to assign all the tasks to $p$, and otherwise, returns that it is a no-instance.
If a branch assigns each task to an index in the plan, it is indeed a yes-instance as the algorithm only assigns executable tasks at each step. For the other direction, assume there is a bijective assignment $a^{*}\colon T\rightarrow[|T|]$ such that, for all $t\in T$, we have $\alpha(t)=p[a^{*}(t)]$ and the corresponding ordering of the tasks according to $a^{*}$ does not conflict with $\prec^{+}$. Consider the branch in which the tasks in $V$ are ordered according to $a^{*}$. As $a^{*}$ is a valid assignment, this branch is not immediately discarded. Further, this branch successfully assigns a task to each index by the following inductive argument. We claim that, for each $0\leq i\leq|T|$, there is a valid assignment that assigns the indices 1 to $i$ exactly like the algorithm does. As $a^{*}$ is a valid assignment, this holds for $i=0$. Suppose the claim holds for an arbitrary but fixed $i$ and the algorithm is about to assign index $i+1$. Then, by the induction hypothesis, there is a valid assignment $a_{i}^{*}$ that assigns all tasks up to $i$ the same as the algorithm has. Let $t$ and $t^{*}$ be the tasks that are assigned to $i+1$ by the algorithm and $a_{i}^{*}$, respectively. If $t=t^{*}$, then the induction claim trivially holds for $i+1$. Thus, assume that $t\neq t^{*}$. As $a_{i}^{*}(t^{*})=i+1$, $t^{*}$ is executable when the algorithm assigns $t$ to $i+1$. Since the algorithm chooses $t$ instead, we have $\psi(t)\leq\psi(t^{*})$. We show that $a_{i+1}^{*}$, the assignment obtained from $a_{i}^{*}$ by switching only $t$ and $t^{*}$, is a valid assignment. Note that the algorithm chooses $t$ at index $i+1$, and so, $t$ is executable at this point. It remains to show that there is no task $t^{\prime}$ such that $a_{i}^{*}(t^{*})<a_{i}^{*}(t^{\prime})<a_{i}^{*}(t)$ and $(t^{*},t^{\prime})\in\prec^{+}$. Observe that when a task from $V$ is executable, it is the only executable task with the lowest priority, and thus, since $\psi(t)\leq\psi(t^{*})$, we get that $t^{*}\notin V$. Hence, as $V$ is a vertex cover, we only have to consider the case where $t^{\prime}\in V$. This implies that if $\psi(t^{*})=\infty$, then no conflicting edge $(t^{*},t^{\prime})$ exists, and so, we only have to consider the case where $\psi(t^{*})\neq\infty$, and thus, $\psi(t)\neq\infty$. Since $a_{i}^{*}$ is a valid assignment, either $t$ is placed before $v_{\psi(t)}$ or $t=v_{\psi(t)-0.5}$. Now, $\psi(t)\leq\psi(t^{*})$ immediately yields that, by swapping $t$ and $t^{*}$, $t^{*}$ is still placed before $v_{\psi(t^{*})}$, and hence, no conflicting edge $(t^{*},t^{\prime})$ exists. Thus, $a_{i+1}^{*}$ is a valid assignment and the induction claim holds for $i+1$. By the inductive argument, the algorithm finds a valid assignment if there is one, and hence, correctly solves the decision problem.
The algorithm branches over the $|V|!$ possible orderings of tasks in the vertex cover. For each ordering, the priorities can be computed in $\mathcal{O}\!\left(|\prec|+|T|\right)$ time. By employing a Hollow Heap Hansen et al. (2017) to implement a priority queue, iterating through the sequence, extracting the minimum priority executable task, and inserting all then executable tasks to the queue takes $\mathcal{O}\!\left(|\prec|+|T|\log|T|\right)$ total time, yielding a runtime of $\mathcal{O}\!\left(|V|!(|\prec|+|T|\log|T|)\right)$ for the entire algorithm. ∎
By Theorem 1, tackling the remaining problems requires additional restrictions on the state space. We show that here, too, we can achieve tractability, albeit the arguments are even more involved than in the previous case.
Theorem 24.
Plan Validation and Action Executability are fixed-parameter tractable when parameterized by the vertex cover number of $G_{\prec}$ plus the number of states in the state transition graph.
Proof.
Let the input instance consist of the primitive network $tn=(T,\prec^{+},\alpha)$, an initial state $s_{0}$, and either a target state $s_{g}$ (for State Reachability) or an action set $S$ (for Action Executability). For Action Executability, assume that Reduction Rule R0 cannot be applied (as otherwise we can solve the instance directly). Begin by computing a minimum-size vertex cover $V$ of $G_{\prec}$ by the classical fixed-parameter algorithm Downey and Fellows (2013), and let $K$ denote the number of states in the state transition graph ${\cal G}$.
The algorithm proceeds analogously for both problems, so we provide only a single description while making explicit the few areas where the techniques diverge. The algorithm branches over all possibilities for the set of employed vertex cover tasks $V^{\prime}\subseteq V$, orders $v_{1},\ldots,v_{|V^{\prime}|}$ of the tasks in $V^{\prime}$, and states $s_{1},\ldots,s_{|V^{\prime}|}$ from which the tasks are executed, respectively. This gives $\mathcal{O}\!\left(2^{|V|}\cdot|V|!\cdot k^{|V|}\right)$ branches.
For each branch, we define a strong equivalence relation on the tasks. For all $t\in T\setminus V$, the admissible interval is defined as $[\max{\{i\mid v_{i}\prec t\}},\min{\{i\mid t\prec v_{i}\}}-1]$. [^2] If the first set is empty, then the interval starts at 0. If the second set is empty, then it ends at $|V^{\prime}|$. Then, $t_{1},t_{2}\in T$ are strongly equivalent if and only if $\alpha(t_{1})$ and $\alpha(t_{2})$ are equivalent and have the same admissible interval. There are at most $E=(k+1)^{k}\cdot(|V^{\prime}|^{2}+1)$ strong equivalence classes. Define the augmented state transition graph ${\cal G}^{\prime}$ over the same set of vertices as the regular state transition graph, but let it have no edges for actions. Instead, let there be an edge $(e,s,s^{\prime})$ for all strong equivalence classes $e$ and states $s,s^{\prime}$ such that $e$ transforms $s$ into $s^{\prime}$. For each such edge, let $\beta((e,s,s^{\prime}))=(s,s^{\prime})$.
Then, split the branch for each of the possible $|V^{\prime}|+1$ simple edge paths in the augmented state transition graph ${\cal G}^{\prime}$ that start and end in the following vertices. A path $W_{0}$ from $s_{0}$ to $s_{1}$, a path $W_{1}$ from $s_{1}\setminus\operatorname{del}(v_{1})\cup\operatorname{add}(v_{1})$ to $s_{2}$, and so on. For Action Executability, also branch on all possibilities for a path $W_{|V^{\prime}|}$ starting at $s_{|V^{\prime}|}\setminus\operatorname{del}(v_{|V^{\prime}|})\cup\operatorname{add}(v_{|V^{\prime}|})$ and ending at any state. For State Reachability, only consider paths $W_{|V^{\prime}|}$ that end in $s_{g}$. If one of these paths cannot be constructed as there is no directed path between the respective vertices, then discard the branch. This splits each branch into ${\mathcal{O}\!\left((2E)^{k(|V^{\prime}|+1)}\right)}$ new branches.
Further, let $C$ be the set of distinct simple edge cycles in ${\cal G}$, where two cycles are considered distinct if and only if they do not contain the exact same vertices, i.e., cycles obtained from other cycles by inverting their directions are not counted as different cycles. Then, $|C|\leq(2E)^{k}$. For each $0\leq i\leq|V^{\prime}|$, we branch over all possible sets $C_{i}\subseteq C$, the cycles that are walked as a detour from path $W_{i}$. This splits each prior branch into ${\mathcal{O}\!\left((2^{|C|})^{|V^{\prime}|+1}\right)}\leq{\mathcal{O}\!\left(2^{(2E)^{k}\cdot(|V^{\prime}|+1)}\right)}$ new branches. Thus, the total number of branches is
$\mathcal{O}\Big{(}2^{|V|}\cdot|V|!\cdot k^{|V|}\cdot(2E)^{k(|V^{\prime}|+1)}\cdot 2^{{(2E)^{k}\cdot(|V^{\prime}|+1)}}\Big{)}.$
Branches where there is a task $v_{i}$ such that $\operatorname{prec}(v_{i})\nsubseteq s_{i}$ are discarded. Further, a branch is discarded if, for some $0\leq i\leq|V^{\prime}|$, there is an equivalence class in $W_{i}$ or in a cycle $c\in C_{i}$ such that $i$ is outside the admissible interval of that class. Branches where there is $0\leq i\leq|V^{\prime}|$ such that $W_{i}$ and its associated cycles do not describe a connected component are discarded as well (as can be tested by the marking procedure described in the proof of Theorem 10).
We build a solution plan that walks each of the paths $W_{i}$, $0\leq i\leq|V^{\prime}|$, and additionally walks through the associated cycles. The exact number of times a cycle is walked is determined by an ILP. This ILP has a variable $x_{c}$ for every $c\in C$. For each strong equivalence class $e$, let $\max{{}^{\prime}}(e)$ denote the number of tasks in $T\setminus V$ in equivalence class $e$. For all $c\in C$, let $c(e)$ denote the number of occurrences of $e$ in $c$. Further, let $w_{i}(e)$ denote the number of occurrences of $e$ in $W_{i}$ for $0\leq i\leq|V^{\prime}|$. Note that, in any plan that is a solution to the problem instance, two strongly equivalent tasks can be swapped and the plan will still be valid. Thus, to ensure no task in $T\setminus V$ is executed more than once, for each equivalence class $e$ we have a constraint
$\max{{}^{\prime}}(e)$ $\geq w_{0}(e)+\dots+w_{|V^{\prime}|}(e)$
$\hphantom{\geq}+c_{1}(e)x_{c_{1}}+\dots+c_{|C|}(e)x_{c_{|C|}},$
where $c_{1},\ldots,c_{|C|}$ are the elements of $C$. Further, to ensure that each selected cycle is executed at least once for each path it is associated with, for all cycles $c\in C$, add a constraint
$x_{c}\geq|{\{0\leq i\leq|V^{\prime}|\mid c\in C_{i}\}}|.$
Only for Action Executability instances, for each action equivalence class $e_{a}$, let $\min^{\prime}(e_{a})$ be the number of occurrences of $e_{a}$ in $S$, and let $w_{i}^{\prime}(e_{a})$ and $c_{i}^{\prime}(e_{a})$ be the number occurrences of $e_{a}$ in path $W_{i}$ and cycle $c_{i}$, respectively. Then, add a constraint
$\min{{}^{\prime}}(e_{a})$ $\geq w_{0}^{\prime}(e_{a})+\dots+w_{|V^{\prime}|}^{\prime}(e_{a})$
$\hphantom{\geq}+c_{1}^{\prime}(e_{a})x_{c_{1}}+\dots+c_{|C|}^{\prime}(e_{a})x_{c_{|C|}}.$
For all instances, the algorithm decides that the given instance is a yes-instance if and only if there is a solution to the ILP for any non-discarded branch.
If the ILP has a solution, then construct a sequence of equivalence classes as follows. For each $0\leq i\leq|V^{\prime}|$, take any cycle $c\in C_{i}$ that shares a vertex $s$ with the current path $W_{i}$. Delete $c$ from $C_{i}$ and augment $W_{i}$ by inserting cycle $c$ at that position once. If there is no more $0\leq j\leq|V|^{\prime}$, $j\neq i$, such that $c\in C_{j}$, instead add as many repetitions of $c$ to $W_{i}$ as the number of times that cycle $c$ was added to the path exactly $x_{c}$ times in total. As the cycles end in the same state they started in, the resulting action sequence is executable. Repeat until $C_{i}$ is empty. As $C_{i}$ and $W_{i}$ form a connected component, all cycles in $C_{i}$ can be added this way. Construct a solution task sequence by concatenating the augmented paths $W_{0},\ldots,W_{|V^{\prime}|}$ and, for a step by a certain equivalence class, employ any task of that class. For Action Executability, choose tasks with a not yet covered action in $S$ whenever possible. Right before each path $W_{i}$,$i\in[|V^{\prime}|]$, execute task $v_{i}$. Observe that this way the preconditions of each task are fulfilled when it is about to be executed. Further, branches in which there are tasks that are executed before their predecessors or after their successors in $\prec^{+}$ are already discarded. The ILP ensures that each equivalence class $e$ appears at most $\max{{}^{\prime}}(e)$ times. As there are $\max{{}^{\prime}}(e)$ tasks of that class and they can be used interchangeably, the sequence is a valid construction. For State Reachability, only paths that end in $s_{g}$ are considered. For Action Executability, the additional set of constraints ensures that $S$ is covered as there are enough tasks of each action and tasks from the same strong equivalence class can be used interchangeably. Thus, the given instance is a yes-instance.
For the other direction, suppose there is a task sequence solving the instance. There is a branch of the algorithm that considers exactly the set $V^{\prime}$ of vertex cover tasks used in the sequence in the order $v_{1},\ldots,v_{|V|}$ and executes each task $v_{i}$ from state $s_{i}$. Split the sequence at the vertex cover tasks into subwalks $W_{0}^{*},\ldots,W_{|V|}^{*}$. Then, for all $0\leq i\leq|V^{\prime}|$, consider an initially empty set of cycles $C^{*}_{i}$. Find any substring $W_{i}^{*}$ that describes a simple edge cycle in ${\cal G}^{\prime}$. Add this cycle to $C^{*}_{i}$ (disregarding direction and starting vertex) and remove the respective substring from $W_{i}^{*}$. Repeat until no more such cycles exist. Observe that each $W_{i}^{*}$ then describes a simple edge path in ${\cal G}^{\prime}$, and all $C^{*}_{i}$ exclusively contain simple edge cycles in ${\cal G}^{\prime}$. We write $W_{i}^{\prime}$ for these reduced $W_{i}^{*}$. There is a branch of the algorithm such that $W_{i}=W_{i}^{\prime}$ and $C_{i}=C_{i}^{*}$ for all $0\leq i\leq|V^{\prime}|$. This branch is not discarded because the task sequence is valid, so $\operatorname{prec}(v_{i})\subseteq s_{i}$ for all $i\in[|V^{\prime}|]$, and the paths $W_{i}$ and cycles in $C_{i}$ are such that they do not depend on any vertex cover task after $v_{i}$, and no vertex cover task up to $v_{i}$ depends on them. Further, as the cycles in $C_{i}^{*}$ were iteratively removed from the path $W_{i}^{*}$, the paths and cycles in $W_{i}$ and $C_{i}$ form a connected component in ${\cal G}^{\prime}$. For each cycle $c\in C$, let $x_{c}^{*}$ be the total number of times $c$ has been removed from paths $W_{i}^{*}$. Then, the ILP has a solution if $x_{c}=x_{c}^{*}$ for all $c\in C$ as follows. As the $x_{c}^{*}$ are computed from a valid sequence, summing over all occurrences of an equivalence class $e$ in paths $W_{i}$ and cycles in $C_{i}$ does not exceed $\max{{}^{\prime}}(e)$. Further, by construction, $x_{c}^{*}$ is at least the number of sets $C_{i}$ with $c\in C_{i}$. In the case of a valid Action Executability solution, the number of occurrences of each action equivalence class $e_{a}$ is at least $\min{{}^{\prime}}(e_{a})$, and thus, the additional constraints are fulfilled. Hence, the ILP has a solution and the algorithm correctly decides for a yes-instance.
Employing reduction rule R0 and building the augmented state transition graph takes $\mathcal{O}\!\left(|A|k+|T|\right)$ time. To decide whether a branch is discarded takes $\mathcal{O}\!\left(|\prec|+|T|+|V||C|^{2}k\right)$ time. Computing $c(e)$, $w_{i}(e)$, $\max^{\prime}(e)$, $c^{\prime}(e_{a})$, $w_{i}^{\prime}(e_{a})$, $\min^{\prime}(e_{a})$, and |{$|{\{0\leq i\leq|V^{\prime}|\mid c\in C_{i}\}}|$ takes $\mathcal{O}\!\left(|T|+(|C|+|V|+1)k+|C|\cdot(|V|+1)\right)$ time. The ILP has $|C|$ variables and at most $E+|C|+(k+1)^{k}={\mathcal{O}\!\left(E+|C|\right)}$ constraints, in which the absolute values of all coefficients are bounded by $\max{\{k,|T|\}}$. As noted before, an ILP instance ${\cal I}_{\text{ILP}}$ with $n$ variables can be solved in $\mathcal{O}\!\left(n^{O(n)}\cdot|{\cal I}_{\text{ILP}}|\right)$ time. Thus, the ILP is solved in
$\mathcal{O}\Big{(}|C|^{O(|C|)}\cdot|C|\cdot(E+|C|)\cdot\log(\max{\{(k,|T|)\}})\Big{)}$
time. As $|C|\leq(2E)^{k}$, the total runtime is at most $f(k,|V|)\cdot|T|$ for some computable function $f$. Multiplying the number of branches with the per-branch runtime yields a quantity that can be upper bounded by $2^{k^{\mathcal{O}(k^{2})}|V|^{\mathcal{O}(k)}}\cdot(|T|^{2})$. ∎
A Meta-Theorem for Fixed-Parameter Tractability.
As our final contribution, we show that the algorithmic meta-theorem for polynomial-time solvability can be lifted to the setting of fixed-parameter tractability if we additionally restrict the maximum “breadth” of a compound task, i.e., ${C_{c}}$.
Theorem 25.
Let PR be a decomposable HTN planning problem and $\kappa$ a stable measure. Assume PR is fixed-parameter tractable on primitive task networks when parameterized by $\kappa$. Then, PR is fixed-parameter tractable when parameterized by $\kappa+{C_{d}}+{C_{\#}}+{C_{s}}+{C_{c}}$.
Proof.
Let ${\cal I}$ have task network $tn=(T,\prec^{+},\alpha)$. If there are no compound tasks in $T$, i.e., ${C_{d}}=0$, both statements immediately hold by applying an algorithm for primitive networks. Thus, assume ${C_{d}}\geq 1$. Then, PR in $tn$ is solved by listing all primitive instances $tn$ can be decomposed into and solving PR in each of these. By Lemma 15, there are at most ${C_{c}}^{\sum_{i=0}^{{C_{d}}-1}{C_{\#}}\cdot{C_{s}}^{i}}$ such instances ${\cal I}_{p}$ and, for each of these, $|{\cal I}_{p}|\leq|{\cal I}|+f^{\prime\prime}({C_{d}},{C_{\#}},{C_{s}})$ for some computable function $f^{\prime\prime}$. Further, for $\kappa^{\prime}$, the parameter $\kappa$ on ${\cal I}_{p}$, we have $\kappa^{\prime}\leq f^{\prime\prime\prime}(\kappa,{C_{d}},{C_{\#}},{C_{s}})$, for some computable function $f^{\prime\prime\prime}$ as $\kappa$ is stable. Thus, if primitive instances ${\cal I}_{p}$ can be solved in time $f(\kappa)\cdot|{\cal I}_{p}|^{\mathcal{O}\!\left(1\right)}$, then the total runtime is
$\mathcal{O}\Big{(}{C_{c}}^{g({C_{d}},{C_{\#}},{C_{s}})}\cdot f(f^{\prime\prime\prime}(\kappa,{C_{d}},{C_{\#}},{C_{s}}))\cdot$
$(|{\cal I}|+f^{\prime\prime}({C_{d}},{C_{\#}},{C_{s}}))^{{\mathcal{O}\!\left(1\right)}}\Big{)}$
$={\mathcal{O}\!\left({C_{c}}^{g({C_{d}},{C_{\#}},{C_{s}})}\cdot f^{\prime}(\kappa,{C_{d}},{C_{\#}},{C_{s}})\cdot|{\cal I}|^{{\mathcal{O}\!\left(1\right)}}\right)},$
where $g({C_{d}},{C_{\#}},{C_{s}})=\sum_{i=0}^{{C_{d}}-1}{C_{\#}}\cdot{C_{s}}^{i}$ and $f^{\prime}(\kappa,{C_{d}},{C_{\#}},{C_{s}})=f(f^{\prime\prime\prime}(\kappa,{C_{d}},{C_{\#}},{C_{s}}))\cdot f^{\prime\prime}({C_{d}},{C_{\#}},{C_{s}})^{\mathcal{O}\!\left(1\right)}$. ∎
7 Concluding Remarks
This article provides a comprehensive understanding of the complexity-theoretic boundaries for several fundamental problems on hierarchical task networks. Our results include stronger algorithmic lower bounds as well as complementary positive results—not only for specific problems of interest, but also in the form of meta-theorems that can be used for any other hierarchical task network planning problem.
In terms of open questions, we note that while we have provided lower bounds justifying all the restrictions applied in the algorithmic meta-theorem for polynomial-time solvability (Theorem 16), in the more refined fixed-parameter tractability setting, we leave it open whether the “hierarchical breadth” ${C_{c}}$ needs to be part of the parameterization in order to establish Theorem 25. While we suspect this to be necessary for at least some of the considered problems, we conclude by showing that it may be omitted in the case of Plan Validation.
Theorem 26.
Let $\kappa$ be a stable measure of HTNs such that Plan Validation is fixed-parameter tractable on primitive networks when parameterized by $\kappa$. Then PR is fixed-parameter tractable parameterized by $\kappa+{C_{d}}+{C_{\#}}+{C_{s}}$.
Proof.
The algorithm employs the same approach as the one used for Theorem 16, but first employs a reduction rule that bounds the number of choices of compound tasks by a function of ${C_{d}},{C_{\#}},{C_{s}}$.
The input plan $p$ gives exactly the multiset of actions that are executed by any solution task sequence. Out of this multiset, exactly $|T|-{C_{\#}}$ actions are covered by primitive tasks in the initial network. Let $S$ denote the multiset of actions that are not covered by primitive tasks. By Lemma 15, any primitive network into which the initial network can be decomposed has at most ${C_{\#}}\cdot{C_{s}}^{{C_{d}}}$ additional tasks. Thus, if $|S|=|p|-(|T|-{C_{\#}})>{C_{\#}}\cdot{C_{s}}^{{C_{d}}}$, the instance is a no-instance. Otherwise, the ${C_{\#}}$ compound tasks in the initial network have to be decomposed to exactly cover $S$. Thus, we can disregard all methods $(c,tn_{m})\in M$ where the primitive tasks in $tn_{m}$ are not a subset of $S$. Out of the remaining methods $(c,tn_{m})\in M$ with non-primitive task network $tn_{m}$, we only require those for which $tn_{m}$ can be decomposed into one of the remaining primitive networks, each of which has depth at most ${C_{d}}-1$. There are at most $2^{|S|}$ possibilities for which actions the primitive networks capture, so by enumerating all possible partial orders, there are less than $2^{|S|}\cdot 2^{|S|^{2}}=2^{|S|^{2}+|S|}$ pairwise non-isomorphic primitive networks left. As we assumed that $|S|\leq{C_{\#}}\cdot{C_{s}}^{{C_{d}}}$, this upper bounds the number of pairwise non-isomorphic primitive networks we have to consider by a function of ${C_{d}},{C_{\#}},{C_{s}}$. Further, the number of methods with pairwise non-isomorphic task networks of depth at most ${C_{d}}-1$ whose primitive tasks are subsets of $S$ and that can be decomposed into one of these at most $2^{|S|^{2}+|S|}$ primitive networks is upper bounded by a function of ${C_{d}},{C_{\#}},{C_{s}}$ as well. This bounds the number of relevant subnetworks in $M$. As there are at most ${C_{\#}}\cdot{C_{s}}^{{C_{d}}}$ compound tasks that are either in the initial network or any relevant subnetwork, the number of relevant compound task names is bounded by a function of ${C_{d}},{C_{\#}},{C_{s}}$ as well. By discarding all methods $(c,tn_{m})$ for which not both $c$ and $tn_{m}$ are relevant, we have ${C_{c}}\leq|M|\leq f({C_{d}},{C_{\#}},{C_{s}})$ for some computable function $f$. Then, employing Theorem 25 gives the desired result. ∎